### v1

# Collision-Free Navigation of Wheeled Mobile Robots: An Integrated Path Planning and Tube-Following Control Approach

Abstract
In this paper, an integrated path planning and tube-following control scheme is proposed for collision-free navigation of a wheeled mobile robot (WMR) in a compact convex workspace cluttered with sufficiently separated spherical obstacles. An analytical path planning algorithm is developed based on Bouligand’s tangent cones and Nagumo’s invariance theorem, which enables the WMR to navigate towards a designated goal location from almost all initial positions in the free space, without entering into augmented obstacle regions with safety margins. We further construct a virtual “safe tube” around the reference trajectory, ensuring that its radius does not exceed the size of the safety margin. Subsequently, a saturated adaptive controller is designed to achieve safe trajectory tracking in the presence of disturbances. It is shown that this tube-following controller guarantees that the WMR tracks the reference trajectory within the predefined tube, while achieving uniform ultimate boundedness of both the position tracking and parameter estimation errors. This indicates that the WMR will not collide with any obstacles along the way. Finally, we report simulation and experimental results to validate the effectiveness of the proposed method.
Index Terms:
Tube-following control, adaptive control, path planning, robot navigation, wheeled mobile robots.
I Introduction
The path planning and control of wheeled mobile robots (WMRs) is an important problem in robotics due to the extensive real-world applications that these systems have, such as cargo transportation, exploration of hazardous environments, automated patrolling [1], to name a few cases. In practical applications, WMRs usually operate in obstacle-cluttered environments, which limits the implementation of traditional motion control methods, thus, motivates the development of collision-free navigation algorithms that can steer WMRs from an initial position to a target goal without colliding with any obstacles along the way [2]. Collision-free robot navigation has garnered significant attention from the robotics and control research communities [3]. Existing solutions to the collision-free navigation problem of robots can be primarily classified into two categories [4]: global (map-based) methods and local (reactive) methods. The former relies on global knowledge of the environment, requiring a priori information about the obstacles (e.g., position and shape). In contrast, the latter utilizes only local knowledge of obstacles in the immediate vicinity of the robot, obtained from on-board sensors. Among the global methods, a simple and computationally efficient solution is the artificial potential field (APF)-based approach [5], which uses a field of virtual potential forces to push the robot towards a target position and pull it away from obstacles; Representative APF-based methods can be found in [6, 7, 8] and references therein. These kind of methods often reach local minima, which hinders achieving global convergence to a designated goal (especially in topologically complex settings). Although navigation functions developed in [9] can help to deal with this problem, they come with extra drawbacks, such as the need for unknown tuning of parameters. Other global methods for collision-free navigation are e.g., heuristic approaches like $A^{\star}$ [10, 11], rapidly exploring random tree (RRT) [12], genetic algorithms [13], etc. As most of them are search-based solutions, their performance can be significantly affected by the scale of the problem [14]. Optimization-based methods [15], while providing alternatives, typically rely on numerical solutions of the constrained optimization problems that result on heavily computational costs. Both the search-based and optimization-based methods are inherently open loop, which makes them vulnerable to disturbances and sensor noise. Thus, the actual robot trajectory may deviates from the desired path, hence, risking collisions with the environment. Local methods that offer reactive solutions for collision-free robot navigation are highly desirable in practice. Particularly in autonomous exploration applications, where WMRs often have limited access to global information and must rely on local sensory data to detect obstacles and perform navigate. Bug algorithms that evolved from maze solving algorithms are one of the simplest reaction navigation approaches for mobile robots in planar environments [16, 17]. Lionis et al. [18] proposed locally computable navigation functions for robotic navigation in unknown sphere worlds. Filippidis and Kyriakopoulos [19] introduced adjustable navigation functions that are capable of autonomously tuning parameters when new obstacles are discovered. Paternain and Ribeiro [20] provided a stochastic extension of these navigation functions with provable convergence properties. However, one caveat is that these reactive function-based methods also suffer from local minima [4]. Arslan and Koditschek [21] proposed a sensor-based reactive method based on a robot-centric spatial decomposition for collision-free navigation with convex obstacles. Huber et al. [22] adopted contraction-based dynamical systems theory to achieve a closed-form solution to the avoidance of moving obstacles. Recently, reinforcement learning methods were used in [23], however, the learning process requires a substantial amount of interaction with the real world. Note that although most of the aforementioned reactive approaches are closed-form, they only consider nominal (holonomic or non-holonomic) robot kinematics, without taking into account disturbances/uncertainties that may arise from external forces, modeling errors, inaccurate controls, sensor noise, data dropout, etc. These unknown actions/terms inevitably degrade the system’s performance, potentially leading to deviations of the desired trajectory, thus, jeopardizing the robot’s safety. In this paper, we study the collision-free navigation problem of differential-drive WMRs operating in a compact convex workspace cluttered with multiple sufficiently separated spherical obstacles. To overcome the non-holonomic constraints inherent in this type of systems, an off-center point is selected as the virtual control point, enabling to transform the original non-holonomic kinematics into a fully-actuated two degrees-of-freedom system [24, 25]. Building upon this model, we propose a novel robot navigation scheme that integrates a path planner and a tube-following controller. By “tube-following” it is meant that the tracking trajectory of the WMR should remain within a virtual safe tube around the reference trajectory to ensure its motion safety. The proposed method can achieve collision-free navigation in cluttered environments from almost all initial configurations in the robot’s free space despite the presence of disturbances. To the best of the authors’ knowledge, this is the first time an integrated path planning and tube-following control scheme is proposed for safe robot navigation in cluttered environments. The effectiveness of the proposed method is validated through numerical simulations and experiments. Specifically, the original contributions of the paper are two-fold:
  1. 1)
    Inspired by [4], a safe path planner is presented based on Bouligand’s tangent cones [26] and Nagumo’s invariance theorem [27], yielding a continuous saturated vector field that can guide the WMR from almost all initial positions in the free space towards a designated goal, without entering into the augmented obstacle regions with safety margins. In essence, the proposed planning algorithm is sensor-based, and the vector field can be generated in real-time using only the local sensing information (i.e., the distance between the WMR and the obstacle boundaries).
  2. 2)
    We propose an adaptive tube-following control method that defines a virtual safe tube around the reference trajectory with its radius no larger than the size of the safety margin. It is shown that the derived controller ensures the tracking of the reference trajectory within the predefined safe tube, while achieving uniform ultimate boundedness of both the trajectory tracking and parameter estimation errors, despite the presence of disturbances. By design, the chosen tube radius ensures that the WMR’s trajectory never enters into the actual obstacle regions, thus achieving the collision-free robot navigation in the perturbed conditions.
[…]
II Problem Statement
II-A Kinematics of wheeled mobile robots
In this paper, we consider a nonholonomic wheeled mobile robot (WMR) with a control point $\bar{\bm{P}}$ located at the midpoint of the axis connecting the two driving wheels, as depicted in Fig. 1. Let $\bar{\mathbf{x}}:=[\bar{x},\bar{y}]^{\top}\in\mathbb{R}^{2}$ be the position vector of $\bar{\bm{P}}$ in the global coordinate frame, and $\bar{\theta}$ be the WMR’s heading angle. The kinematic model of the WMR is expressed as
$\left[\begin{matrix}\dot{\bar{x}}\\ \dot{\bar{y}}\\ \dot{\bar{\theta}}\end{matrix}\right]=\left[\begin{matrix}\cos\bar{\theta}&0\\ \sin\bar{\theta}&0\\ 0&1\end{matrix}\right](\mathbf{u}+\mathbf{u}_{d}),$ (1)
where $\mathbf{u}:=[v,\omega]^{\top}\in\mathbb{R}^{2}$ is the the control input vector with $v$ and $\omega$ being the linear and angular velocities of the WMR, respectively, and $\mathbf{u}_{d}\in\mathbb{R}^{2}$ is the matched disturbance that may arise from control errors, sensor noise, etc.
Figure 1: Sketch of the wheeled mobile robot.
To facilitate the subsequent design and analysis, we select an off-center point $\bm{P}$ as the virtual control point, which locates at a distance $\ell\neq 0$ away from $\bar{\bm{P}}$ along the longitudinal axis of the WMR [24, 25], as shown in Fig. 1. Denote by $\mathbf{x}:=[x,y]^{\top}\in\mathbb{R}^{2}$ and $\theta$ the position and heading angle of $\bm{P}$, respectively. We have the following change of coordinates:
[…]
In view of (1) and (2), the kinematics of $\mathbf{x}$ is given by
$\dot{\mathbf{x}}=\mathbf{R}(\theta)(\mathbf{u}+\mathbf{u}_{d}),$ (3)
where $\mathbf{R}(\theta)\in\mathbb{R}^{2\times 2}$ is defined as
$\mathbf{R}(\theta):=\left[\begin{matrix}\cos\theta&-\ell\sin\theta\\ \sin\theta&\ell\cos\theta\\ \end{matrix}\right],$ (4)
which is full rank. By setting $|\ell|\leq 1$, it follows that $\|\mathbf{R}(\theta)\|=\max\{1,|\ell|\}=1$ and $\|\mathbf{R}^{-1}(\theta)\|=\max\{1,1/|\ell|\}=1/\ell$.
Assumption 1.
The disturbance $\mathbf{u}_{d}$ is bounded by an unknown constant $d=\sup_{t\geq 0}\|\mathbf{u}_{d}(t)\|$, that is, $\|\mathbf{u}_{d}\|\leq d$.
Remark 1.
The full rank of $\mathbf{R}(\theta)$ enables us to freely control the position of the WMR (specifically, the virtual control point $\bm{P}$) without being restricted by nonholonomic constraints.
II-B Operating environment
Consider a WMR operating inside a closed compact convex workspace $\mathcal{W}^{\ast}\subset\mathbb{R}^{2}$, punctured by a set of $n\in\mathbb{N}$ obstacles $\mathcal{O}_{i}^{\ast}$, $i\in\mathbb{I}:=\{1,2,...,n\}$, which are represented by open balls with centers $\mathbf{c}_{i}\in\mathbb{R}^{2}$ and radii $r_{i}>0$. A circumscribed circle centered at $\bm{P}$ with radius $r>0$ is constructed, which is the smallest circle enclosing the WMR (see Fig. 1). To ensure that the WMR can navigate freely between any of the obstacles in $\mathcal{W}^{\ast}$, we make the following assumption:
Assumption 2.
The $n$ obstacles are separated from each other by a clearance of at least
${\rm d}(\mathcal{O}_{i}^{\ast},\mathcal{O}_{j}^{\ast})>2(r+h),~{}\forall i,j\in\mathbb{I},~{}i\neq j,$ (5)
and from the boundary of the workspace $\mathcal{W}^{\ast}$ as
${\rm d}(\mathcal{O}_{i}^{\ast},\partial\mathcal{W}^{\ast})>2r+h,~{}\forall i\in\mathbb{I},$ (6)
where $h>0$ is a constant.
For ease of design, the WMR is considered as a point (i.e., the virtual control point $\bm{P}$), by transferring the volume of the circumscribed circle to the other workspace entities. Then, for the point $\bm{P}$, the workspace is defined as
$\mathcal{W}:=\{\mathbf{x}\in\mathbb{R}^{2}\mid\text{d}_{\complement\mathcal{W}^{\ast}}(\mathbf{x})\geq r\},$ (7)
and the obstacle regions are defined as the following pairwise disjoint spherical subsets of $\mathbb{R}^{2}$:
$\mathcal{O}_{i}:=\{\mathbf{x}\in\mathbb{R}^{2}\mid\beta_{i}(\mathbf{x})<0\},~{}i\in\mathbb{I},$ (8)
where $\beta_{i}(\mathbf{x}):=\|\mathbf{x}-\mathbf{c}_{i}\|-(r+r_{i})$. Thus, the free space of $\bm{P}$ is given by the closed set $\mathcal{X}:=\mathcal{W}\setminus\mathcal{O}$ with $\mathcal{O}:=\cup_{i=1}^{n}\mathcal{O}_{i}$. To enhance the safety, a safety margin of size $0<\epsilon<h$ is introduced (see the light blue regions in Fig. 2), which results in an eroded workspace $\mathcal{W}^{\epsilon}:=\{\mathbf{x}\in\mathbb{R}^{2}\mid\text{d}_{\complement\mathcal{W}^{\ast}}(\mathbf{x})\geq r+\epsilon\}$ and $n$ augmented obstacle regions $\mathcal{O}_{i}^{\epsilon}:=\{\mathbf{x}\in\mathbb{R}^{2}\mid\beta_{i}(\mathbf{x})<\epsilon\}$, $i\in\mathbb{I}$. By doing so, the free space of $\bm{P}$ reduces to $\mathcal{X}_{\epsilon}:=\mathcal{W}^{\epsilon}\setminus\mathcal{O}^{\epsilon}$ with $\mathcal{O}^{\epsilon}:=\cup_{i=1}^{n}\mathcal{O}_{i}^{\epsilon}$.
Figure 2: Schematic diagram of the operating environment. The dark gray balls denote the obstacles $\mathcal{O}_{i}^{\ast}$, $i\in\mathbb{I}$, the light gray regions denote the augmented workspace boundary and obstacles, whereas the light blue regions denote the safety margin.
II-C Problem formulation
The collision-free navigation problem for the WMR is now formulated as follows:
Problem 1.
[…]
III Main Results
To solve Problem 1, we propose a closed-form robot navigation scheme that integrates a path planner and a tube-following controller. The block diagram of this scheme is depicted in Fig. 3. The path planner explicitly considers the safety margin and generates a collision-free reference trajectory $\mathbf{x}_{d}$ in the free space $\mathcal{X}{\epsilon}$, using tangent cones and the well-known Nagumo’s invariance theorem. In the control module, an adaptive tube-following controller is designed to track the planned trajectory in the presence of unknown disturbances, while ensuring that the position $\mathbf{x}$ of $\bm{P}$ stays within a predefined tube around the reference trajectory $\mathbf{x}_{d}$. Under this framework, if the radius of the tube is set less than the width $\epsilon$ of the safety margin, then the actual motion trajectory of $\mathbf{x}$ will remain within the set $\mathcal{X}$, thus achieving safe robot motions.
Figure 3: Block diagram of the collision-free navigation scheme.
III-A Preliminaries
The definition of tangent cones and Nagumo’s theorem that are necessary for collision-free path planning are given.
Definition 1 (Bouligand’s tangent cone [26]).
Given a closed set $\mathcal{F}\in\mathbb{R}^{n}$, the tangent cone to $\mathcal{F}$ at a point $\mathbf{x}\in\mathbb{R}^{n}$ is the subset of $\mathbb{R}^{n}$ defined by
$\mathbf{T}_{\mathcal{F}}(\mathbf{x}):=\left\{\mathbf{z}\in\mathbb{R}^{n}\mid\mathop{\lim\inf}\limits_{\tau\to 0^{+}}\dfrac{{\rm d}_{\mathcal{F}}(\mathbf{x}+\tau\mathbf{z})}{\tau}=0\right\}.$
The tangent cone $\mathbf{T}_{\mathcal{F}}(\mathbf{x})$ is a set that contains all the vectors pointing from $\mathbf{x}$ inside or tangent to $\mathcal{F}$, while for $\mathbf{x}\notin\mathcal{F}$, $\mathbf{T}_{\mathcal{F}}(\mathbf{x})=\emptyset$. Since for all $\mathbf{x}\in\text{int}(\mathcal{F})$, we have $\mathbf{T}_{\mathcal{F}}(\mathbf{x})=\mathbb{R}^{n}$, the tangent cone $\mathbf{T}_{\mathcal{F}}(\mathbf{x})$ is non-trivial only on the boundary $\partial\mathcal{F}$. Next, we recall the Nagumo’s invariance theorem.
Theorem 1 (Nagumo 1942 [27]).
Consider the system $\dot{\mathbf{x}}(t)=\mathbf{f}(\mathbf{x}(t))$ and assume that it admits a unique solution in forward time for each initial condition $\mathbf{x}(0)$ in an open set $\mathcal{O}$. Then, the closed set $\mathcal{F}\subset\mathcal{O}$ is forward invariant iff
$\mathbf{f}(\mathbf{x})\in\mathbf{T}_{\mathcal{F}}(\mathbf{x}),~{}\forall\mathbf{x}\in\mathcal{F}.$ (9)
An intuitive interpretation of Nagumo’s theorem is that if the vector field $\mathbf{f}(\mathbf{x})$ points inside or is tangent to the set $\mathcal{F}$ at each point $\mathbf{x}\in\mathcal{F}$, then $\mathcal{F}$ is forward invariant. As stated earlier, $\mathbf{T}_{\mathcal{F}}(\mathbf{x})$ is non-trivial only on $\partial\mathcal{F}$, thus it is only necessary to check (9) for the boundary points.
Lemma 1.
Given any constant $b>0$, the following inequalities hold for all $0\leq q<b$:
$\frac{q}{b}\leq\ln\dfrac{b}{b-q}\leq\dfrac{q}{b-q}.$
Proof. Define two functions $h_{1}(q):=q/(b-q)-\ln(b/(b-q))$ and $h_{2}(q):=\ln(b/(b-q))-q/b$. After simple calculations, it follows that
$\dfrac{\partial h_{1}(q)}{\partial q}=\dfrac{q}{(b-q)^{2}}\geq 0,~{}\dfrac{\partial h_{2}(q)}{\partial q}=\dfrac{q}{b(b-q)}\geq 0,$
[…]
III-B Path planing based on tangent cones
Based on Bouligand’s tangent cones and Nagumo’s theorem, a safe path planning approach is proposed, which can guide the virtual control point $\bm{P}$ towards the goal $\mathbf{x}^{\ast}$ (motion-to-goal) from almost [^1] The free space $\mathcal{X}_{\epsilon}$ is a non-contractible space, for which the topological obstruction precludes the possibility of any continuous vector fields achieving global asymptotic stability. The basin of attraction of the desired equilibrium must generally exclude at least a set of measure zero. all initial conditions $\mathbf{x}(0)\in\mathcal{X}_{\epsilon}$, while ensuring forward invariance of the free space $\mathcal{X}_{\epsilon}$ (collision avoidance). We neglect the disturbance $\mathbf{u}_{d}$ in (3) and consider a control law of the following form
$\mathbf{u}=\mathbf{R}^{-1}(\theta)\bm{\tau},$ (10)
where $\bm{\tau}\in\mathbb{R}^{2}$ is a virtual control law to be designed. Then, substituting (10) into (3), one gets
$\dot{\mathbf{x}}=\bm{\tau}.$ (11)
According to Nagumo’s theorem, it is sufficient to address the collision-free path planning problem by solving the following constrained optimization problem
$\begin{split}\min\limits_{\bm{\tau}}&~{}\|\bm{\tau}-\bm{\kappa}_{0}(\mathbf{x})\|,\\ \text{s.t.}&~{}\bm{\tau}\in\mathbf{T}_{\mathcal{{X}}_{\epsilon}}(\mathbf{x}),~{}\forall\mathbf{x}\in\mathcal{X}_{\epsilon},\end{split}$ (12)
where $\bm{\kappa}_{0}(\mathbf{x})$ is a nominal control law for motion-to-goal and is designed here as the following smooth function:
$\bm{\kappa}_{0}(\mathbf{x})=-k_{0}(\mathbf{x})(\mathbf{x}-\mathbf{x}^{\ast}),$ (13)
with $k_{0}(\mathbf{x})>0$ given by
$k_{0}(\mathbf{x}):=\dfrac{\alpha}{\sqrt{\|\mathbf{x}-\mathbf{x}^{\ast}\|^{2}+\beta^{2}}},$ (14)
where $\alpha$ and $\beta$ are positive design constants. From (13) and (14), it follows that $\|\bm{\kappa}_{0}(\mathbf{x})\|\leq\alpha$, which allows us to generate a saturated nominal vector field. The problem (12) is clearly a nearest point problem, which has an explicit solution that will be demonstrated in the following.
Figure 4: The resulting smooth navigation trajectories maintains a safety margin of size $\epsilon$ (light blue regions) from the workspace boundary and obstacles. The dashed circles denote the influence regions from each obstacle. Moreover, the orange points are the unstable stationary points associated with each obstacles.
The robot workspace $\mathcal{W}^{\ast}$ is a convex set, so does $\mathcal{W}^{\epsilon}$, which together with the fact that $\mathbf{x}^{\ast}\in\text{int}(\mathcal{X}_{\epsilon})$ suggests that for all $\mathbf{x}\in\partial\mathcal{W}^{\epsilon}$, $\bm{\kappa}_{0}(\mathbf{x})$ points inside the free space $\mathcal{X}_{\epsilon}$ (i.e., $\bm{\kappa}_{0}(\mathbf{x})\in\mathbf{T}_{\mathcal{{X}}_{\epsilon}}(\mathbf{x})$) and thus is a solution to (12). Moreover, for all $\mathbf{x}\in\text{int}(\mathcal{X}_{\epsilon})$, it follows that $\bm{\kappa}_{0}(\mathbf{x})$ is also the solution to (12). Next we check the obstacle boundary points $\mathbf{x}\in\partial\mathcal{O}^{\epsilon}$. As $\partial\mathcal{O}^{\epsilon}$ is a smooth hypersurface of $\mathbb{R}^{2}$, which is orientable, there exists a continuously differentiable map (known as the Gauss map [28]) $\mathbf{n}:\partial\mathcal{O}^{\epsilon}\to\mathbb{S}^{1}$ such that for all $\mathbf{x}\in\partial\mathcal{O}^{\epsilon}$, $\mathbf{n}(\mathbf{x})$ is the outward unit normal vector to $\partial\mathcal{O}^{\epsilon}$. As clearly seen in Fig. 2, the tangent cone at any $\mathbf{x}\in\partial\mathcal{O}^{\epsilon}$ is a half-space
$\mathbf{T}_{\mathcal{X}_{\epsilon}}(\mathbf{x}):=\{\mathbf{y}\in\mathbb{R}^{2}\mid(\mathbf{y}-\mathbf{x})^{\top}\mathbf{n}(\mathbf{x})\leq 0\},$ (15)
which is a convex function. Thus, (12) has a unique solution. Based on the above argument, it is known that if $\mathbf{x}\in\mathcal{X}_{\epsilon}\setminus\partial\mathcal{O}^{\epsilon}$ or $\mathbf{x}\in\partial\mathcal{O}^{\epsilon}$ and $\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{n}(\mathbf{x})\leq 0$, then $\bm{\kappa}_{0}(\mathbf{x})$ is a solution to (12). On the other hand, if $\mathbf{x}\in\mathcal{O}^{\epsilon}$ and $\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{n}(\mathbf{x})>0$, the closest point of (12) is obtained by the orthogonal projection $\mathbf{Q}(\mathbf{n}(\mathbf{x})):=\mathbf{I}_{2}-\mathbf{n}(\mathbf{x})\mathbf{n}^{\top}(\mathbf{x})$ onto the tangent hyperplane of $\partial\mathcal{X}_{\epsilon}$, defined by $\mathbf{H}_{\partial\mathcal{X}_{\epsilon}}(\mathbf{x}):=\{\mathbf{y}\in\mathbb{R}^{2}\mid(\mathbf{y}-\mathbf{x})^{\top}\mathbf{n}(\mathbf{x})=0\}$. Then, a general solution to (12) is as follows:
$\bm{\tau}(\mathbf{x})=\left\{\begin{split}&\bm{\kappa}_{0}(\mathbf{x}),&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})>\epsilon~{}\text{or}\\ &&&\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{n}(\mathbf{x})\leq 0,\\ &\mathbf{Q}(\mathbf{n}(\mathbf{x}))\bm{\kappa}_{0}(\mathbf{x}),&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})=\epsilon~{}\text{and}\\ &&&\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{n}(\mathbf{x})>0.\end{split}\right.$ (16)
By inspecting (16), we find that the resulting vector field is discontinuous at some boundary points $\mathbf{x}\in\partial\mathcal{O}^{\epsilon}$ like $\bm{P}_{1}-\bm{P}_{4}$ in Fig. 2. To address this issue, a continuous version of (16) is proposed in the sequel. Following the line of [4], we specify an influence region for each obstacle (marked by the dashed line in Fig. 4), which is defined as $\{\mathbf{x}\in\mathbb{R}^{2}\mid\text{d}_{\mathcal{O}_{i}}(\mathbf{x})\leq\epsilon^{\ast}\}$, where $\epsilon^{\ast}\in(\epsilon,h]$ and $i\in\mathbb{I}$. Obstacle avoidance is activated only when the position $\mathbf{x}$ of $\bm{P}$ enters the influence regions of obstacles. To proceed, a bearing vector is defined as
$\mathbf{b}(\mathbf{x}):=\dfrac{\mathbf{P}_{\partial\mathcal{O}}(\mathbf{x})-\mathbf{x}}{\|\mathbf{P}_{\partial\mathcal{O}}(\mathbf{x})-\mathbf{x}\|}=\dfrac{\mathbf{P}_{\partial\mathcal{O}}(\mathbf{x})-\mathbf{x}}{\text{d}_{\mathcal{O}}(\mathbf{x})}.$ (17)
for all $\mathbf{x}$ with $\epsilon\leq\text{d}_{\mathcal{O}}(\mathbf{x})\leq\epsilon^{\ast}$. Since $\text{d}(\mathcal{O}_{i}^{\ast},\mathcal{O}_{j}^{\ast})>2(r+h)$, $\forall i,j\in\mathbb{I}$, $i\neq j$ (referring to (5) in Assumption 2) and $\epsilon^{\ast}\leq h$, there can be only one obstacle $\mathcal{O}_{i}$, $i\in\mathbb{I}$ such that $\text{d}_{\mathcal{O}}(\mathbf{x})=\text{d}_{\mathcal{O}_{i}}(\mathbf{x})\leq\epsilon^{\ast}$. With this in mind, $\mathbf{b}(\mathbf{x})$ in (17) can be computed by $\mathbf{b}(\mathbf{x})=(\mathbf{c}_{i}-\mathbf{x})/\|\mathbf{c}_{i}-\mathbf{x}\|$. Note that when $\mathbf{x}\in\partial\mathcal{O}_{\epsilon}$, $\mathbf{b}(\mathbf{x})$ is equivalent to the Gauss map $\mathbf{n}(\mathbf{x})$. Now the control law (16) is modified as
$\bm{\tau}(\mathbf{x})=\left\{\begin{split}&\bm{\kappa}_{0}(\mathbf{x}),&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})>\epsilon^{\ast}~{}\text{or}\\ &&&\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})\leq 0,\\ &\mathbf{\Pi}(\mathbf{x})\bm{\kappa}_{0}(\mathbf{x}),&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})\leq\epsilon^{\ast}~{}\text{and}\\ &&&\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})>0,\end{split}\right.$ (18)
with
$\mathbf{\Pi}(\mathbf{x}):=\mathbf{I}_{2}-\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))\mathbf{b}(\mathbf{x})\mathbf{b}^{\top}(\mathbf{x}),$ (19)
where $\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))\in\mathbb{R}$ is a $C^{1}$ bump function that smoothly transitions from $1$ to 00 on the interval $\text{d}_{\mathcal{O}}(\mathbf{x})\in[\epsilon,\epsilon^{\ast}]$, that is to say, its first derivatives equal to zero at the endpoints of the interval $[\epsilon,\epsilon^{\ast}]$. A simple choice of $\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))$ is
$\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))=\left\{\begin{split}&1,&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})\leq\epsilon,\\ &\lambda(\text{d}_{\mathcal{O}}(\mathbf{x})),&\text{if}~{}&\epsilon<\text{d}_{\mathcal{O}}(\mathbf{x})<\epsilon^{\ast},\\ &0,&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})\geq\epsilon^{\ast},\end{split}\right.$ (20)
with
$\lambda(\text{d}_{\mathcal{O}}(\mathbf{x}))=\dfrac{1}{2}\left[1-\cos\left(\pi\dfrac{\epsilon^{\ast}-\text{d}_{\mathcal{O}}(\mathbf{x})}{\epsilon^{\ast}-\epsilon}\right)\right].$ (21)
One can verify that $\phi(\epsilon)=1$, $\phi(\epsilon^{\ast})=0$, and
$\left.\dfrac{d\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))}{d\text{d}_{\mathcal{O}}(\mathbf{x})}\right|_{\text{d}_{\mathcal{O}}(\mathbf{x})=0~{}\text{or}~{}\epsilon^{\ast}}=0.$
Since the nominal control law, $\bm{\kappa}_{0}(\mathbf{x})$, the bearing vector, $\mathbf{b}(\mathbf{x})$, and the bump function, $\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))$, are continuously differentiable, the resulting control law $\bm{\tau}(\mathbf{x})$ in (18) is continuous and piecewise continuously differentiable on the domain $\mathcal{X}$. It is not differentiable only at those points satisfying $\text{d}_{\mathcal{O}}(\mathbf{x})\leq\epsilon^{\ast}$ and $\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})=0$. Now the main result of the path planning problem is summarized in the following theorem.
Theorem 2.
[…]
  1. 1.
    The free space $\mathcal{X}_{\epsilon}$ in (7) is forward invariant.
  2. 2.
    For any $\mathbf{x}(0)\in\mathcal{X}_{\epsilon}$, the kinematics (11) admits a unique solution in forward time, which asymptotically converges to the set $\{\mathbf{x}^{\ast}\}\cup_{i=1}^{n}\mathcal{S}_{i}$ with $\mathcal{S}_{i}$ given by
    $\mathcal{S}_{i}:=\left\{\mathbf{x}\in\mathcal{X}_{\epsilon}\bigg{|}{\rm d}_{\mathcal{O}_{i}}(\mathbf{x})=\epsilon,~{}\dfrac{\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})}{\|\bm{\kappa}_{0}(\mathbf{x})\|}=1\right\}.$ (22)
  3. 3.
    The undesired stationary points $\mathbf{s}_{i}\in\mathcal{S}_{i}$, associated with obstacles $\mathcal{O}_{i}$, are locally unstable and only have a stable manifold of measure zero.
  4. 4.
    The equilibrium point $\mathbf{x}=\mathbf{x}^{\ast}$ is almost globally asymptotically stable and locally exponentially stable.
[…]
Remark 2.
The proposed safe path planning scheme offers an analytical solution and thus is computationally much more efficient than the optimization-based methods. Moreover, unlike the navigation functions, the scheme can achieve almost global asymptotic convergence without requiring unknown parameter tuning, making it easy to implement in practice.
Remark 3.
A smooth and saturated nominal control law (18) is proposed, instead of the traditional proportional control law $k_{0}(\mathbf{x})=-k(\mathbf{x}-\mathbf{x}^{\ast})$ [4, 21], with the aim of generating a saturated vector field (18). This is justified by the inequality $\|\bm{\tau}(\mathbf{x})\|\leq\|\bm{\kappa}_{0}(\mathbf{x})\|\leq\alpha$, where the equality $\|\bm{\Pi}(\mathbf{x})\|=1$ has been used. In fact, different types of saturation functions can be used to design $\kappa_{0}(\mathbf{x})$ as an alternative, such as the standard $\text{sat}(\cdot)$, hyperbolic tangent function $\tanh(\cdot)$, $\text{arctan}(\cdot)$, etc. However, it is worth noting that these functions may result in the nominal control direction not aligning directly with the desired position, which in turn leads to an unnecessarily long path.
III-C Adaptive tube-following control
The planning algorithm introduced in Sec. III-B can generate a collision-free reference trajectory for the WMR. To avoid conflicts of notations, the reference trajectory is denoted as $\mathbf{x}_{d}$, which is governed by
$\dot{\mathbf{x}}_{d}=\bm{\tau}_{d},$ (23)
where $\bm{\tau}_{d}$ is given by (18) but with $\mathbf{x}$ replaced by $\mathbf{x}_{d}$. Let us define the position tracking error as $\mathbf{x}_{e}:=\mathbf{x}-\mathbf{x}_{d}$ and impose an inequality constraint on it as follows:
$\|\mathbf{x}_{e}(t)\|^{2}<\rho^{2},$ (24)
where $\rho>0$ is a user-defined constant. At each time instance, the actual position $\mathbf{x}$ of the virtual control point $\bm{P}$ is allowed to stay within a 1-sphere of radius $\rho$ and center $\mathbf{x}_{d}(t)$. Unifying all such spheres along $t$ forms a tube around $\mathbf{x}_{d}(t)$, as shown in Fig. 5. Since a safety margin is considered in the planning module, if the tube radius is taken no larger than the size $\epsilon$ of the safety margin and (24) holds for all $t\geq 0$, then the actual trajectory $\mathbf{x}(t)$ will remain within the predefined tube without colliding with any actual obstacles, that is, $\mathbf{x}\in\mathcal{X}$.
Figure 5: Illustration of the predefined tube.
In the following, an adaptive tube-following controller is developed to achieve high-precision trajectory tracking, while ensuring that $\mathbf{x}(t)$ evolves within the tube defined by (24). To this end, we define a transformed error
$\xi(t):=\dfrac{\|\mathbf{x}_{e}(t)\|^{2}}{\rho^{2}}.$ (25)
Taking the time derivative of (25), whilst using (3) and (23), one can easily get
$\dot{\xi}=\dfrac{2}{\rho^{2}}\mathbf{x}_{e}^{\top}[\mathbf{R}(\theta)(\mathbf{u}+\mathbf{u}_{d})-\bm{\tau}_{d}].$ (26)
Inspecting (25) reveals that $10\leq\xi(t)<$ 1 is equivalent to (24), and $\xi(t)=0$ only when $\mathbf{x}_{e}(t)=\mathbf{0}$. Therefore, the tube-following control problem boils down to achieving $\lim_{t\to\infty}\xi(t)=0$, while ensuring $10\leq\xi(t)<$ 1, $\forall t\geq 0$, via a properly-designed controller. To achieve this, the following logarithmic barrier function is considered
$L=\dfrac{1}{2}\ln\dfrac{1}{1-\xi}.$ (27)
Let $\mathbf{z}:=\mathbf{x}_{e}/(\rho^{2}(1-\xi))$. Then, evaluating the time derivative of $L$ along (26) yields
$\dot{L}=$ $~{}\mathbf{z}^{\top}[\mathbf{R}(\theta)(\mathbf{u}+\mathbf{u}_{d})-\bm{\tau}_{d}]$
$\leq$ $~{}\mathbf{z}^{\top}(\mathbf{R}(\theta)\mathbf{u}-\bm{\tau}_{d})+d\|\mathbf{z}\|,$ (28)
wherein we have used the fact that $\|\mathbf{R}(\theta)\|=1$. The adaptive control law is designed as
$\mathbf{u}=\mathbf{R}^{-1}(\theta)(-k\mathbf{x}_{e}+\bm{\tau}_{d}-\bm{\varpi}),$ (29)
with
$\bm{\varpi}:=\dfrac{\hat{d}^{2}\mathbf{z}}{\sqrt{\hat{d}^{2}\|\mathbf{z}\|^{2}+\varphi^{2}}},$ (30)
where $\hat{d}$ denotes the estimate of $d$ (see Assumption 1 for its definition) and is updated by a projection-based adaptive law. Suppose $d\in\Omega:=\{\nu\in\mathbb{R}\mid 0\leq\nu\leq d_{m}\}$ and let $\Omega_{\delta}:=\{\nu\in\mathbb{R}\mid 0\leq\nu\leq d_{m}+\delta\}$, where $d_{m}>0$ and $\delta>0$ are known constants. Inspired by [29], the adaptive law is derived as
$\dot{\hat{d}}=\text{Proj}(\hat{d},\Phi),~{}\Phi:=\|\mathbf{z}\|-\gamma\hat{d},$ (31)
with
$\text{Proj}(\hat{d},\Phi)=\left\{\begin{split}&\eta\Phi,&{\rm if}~{}&1)~{}0\leq\hat{d}<d_{m}~{}\text{or}\\ &&~{}&2)~{}\hat{d}\geq d_{m}~{}\text{and}~{}\Phi\leq 0,\\ &\eta\breve{\Phi},&{\rm if}~{}&3)~{}\hat{d}\geq d_{m}~{}\text{and}~{}\Phi>0,\end{split}\right.$ (32)
where $\eta,\,\gamma>0$ are design constants and
$\breve{\Phi}=\left(1-\dfrac{\hat{d}-d_{m}}{\delta}\right)\Phi.$ (33)
Note that the projection operator $\text{Proj}(\hat{d},\Phi)$ is locally Lipschitz in $(\hat{d},\Phi)$ [29]. Choosing the initial condition of (31) as $\hat{d}(0)\in\Omega_{\delta}$, we will now demonstrate that
$\hat{d}(0)\in\Omega_{\delta}\Rightarrow\hat{d}(t)\in\Omega_{\delta},~{}\forall t\geq 0.$ (34)
In Cases 1 and 2, the adaptive law (31) reduces to
$\dot{\hat{d}}=\eta(\|\mathbf{z}\|-\gamma\hat{d}).$ (35)
Solving (35) yields
[…]
where $t_{0}\geq 0$ marks the beginning of each time interval during which Cases 1 and 2 hold. Moreover, in these two cases, (34) holds as a result of the given conditions. In Case 3, however, (31) becomes
$\dot{\hat{d}}=\dfrac{\delta+d_{m}-\hat{d}}{\delta}\Phi,$ (37)
which implies that $\hat{d}$ decreases when $\hat{d}>\delta+d_{m}$ and increases if $\hat{d}<\delta+d_{m}$ until it reaches $\delta+d_{m}$ such that $\dot{\hat{d}}=0$. Therefore, the adaptive law (31) ensures that (34) holds. Let us define the parameter estimation error as $\tilde{d}:=\hat{d}-d$.
Theorem 3.
Consider the WMR kinematics (3) under Assumption 1 and suppose that $d\in\Omega$. Given the initial position $\mathbf{x}(0)$ satisfying (25), then the adaptive controller (29), together with the projection-based update law (31), can steer $\mathbf{x}$ to track the reference trajectory $\mathbf{x}_{d}$ given by (23) along with the predefined tube (25), while ensuring uniform ultimate boundedness of the position error $\mathbf{x}_{e}$ and the estimation error $\tilde{d}$.
Proof. Consider the Lyapunov function candidate (LFC):
$V=L+\dfrac{1}{2\eta}\tilde{d}^{2}.$ (38)
Taking the derivative of $V$ and using (III-C) and (29) lead to
$\dot{V}\leq-k\mathbf{z}^{\top}\mathbf{x}_{e}-\mathbf{z}^{\top}\bm{\varpi}+\hat{d}\|\mathbf{z}\|+\dfrac{1}{\eta}\tilde{d}(\dot{\hat{d}}-\eta\|\mathbf{z}\|).$ (39)
Invoking the inequality $|\nu|\leq\nu^{2}/\sqrt{\nu^{2}+\varphi^{2}}+\varphi$ for any $\nu\in\mathbb{R}$ and $\varphi>0$, one gets $-\mathbf{z}^{\top}\bm{\varpi}+\hat{d}\|\mathbf{z}\|\leq\varphi$. In addition, noting (26), (III-C), and Lemma 1, it follows that $-k\mathbf{z}^{\top}\mathbf{x}_{e}=-k\xi/(1-\xi)\leq-2kL$ for all $10\leq\xi<$ 1. As a result, (39) reduces to
$\dot{V}\leq-2kL+\varphi+\dfrac{1}{\eta}\tilde{d}(\dot{\hat{d}}-\eta\|\mathbf{z}\|),$ (40)
on the set $10\leq\xi<$ 1. To proceed, we consider three cases in (32). For Cases 1 and 2, it follows from (35) that
$\tilde{d}(\dot{\hat{d}}-\eta\|\mathbf{z}\|)=-\eta\gamma\tilde{d}\hat{d}.$ (41)
While for Case 3, we have
$\tilde{d}(\dot{\hat{d}}-\eta\|\mathbf{z}\|)=-\eta\gamma\tilde{d}\hat{d}-\dfrac{\eta\tilde{d}(\hat{d}-d_{m})\Phi}{\delta}\leq-\eta\gamma\tilde{d}\hat{d},$ (42)
which is true because $\tilde{d}\geq 0$, $\hat{d}-d_{m}\geq 0$, and $\Phi>0$ according to the conditions specified in Case 3. In view of (41) and (42), one can easily get
$\dot{V}$ $\leq-2kL-\gamma\tilde{d}\hat{d}+\varphi$
$\leq-2kL-\dfrac{\gamma}{2}\tilde{d}^{2}+c$
$\leq-aV+c,$ (43)
where $a:=\min\{2k,\eta\gamma\}>0$ and $c:=\gamma d^{2}/2+\varphi>0$. Integrating both sides of (III-C) over $[0,\infty)$ yields
$V(t)\leq\left(V(0)-\dfrac{c}{a}\right)e^{-at}+\dfrac{c}{a},~{}\forall t\geq 0,$ (44)
showing that $V\in\mathcal{L}_{\infty}$ and hence $L\in\mathcal{L}_{\infty}$. The later implies that $10\leq\xi(t)<$ 1 for all $t\geq 0$. As a consequence, $\mathbf{x}$ keeps within the predefined tube (25). As $L(t)\leq V(t)$, it follows from (25) and (27) that $\|\mathbf{x}_{e}(t)\|\leq\rho\sqrt{1-e^{-2V(t)}}$. In addition, $\tilde{d}^{2}/(2\eta)\leq V(t)$ leads to $|\tilde{d}(t)|\leq\sqrt{2\eta V(t)}$. As shown in (44), $V(t)$ is eventually bounded by $c/a$, whereby one can get
$\mathop{\lim\sup}\limits_{t\to\infty}\|\mathbf{x}_{e}(t)\|\leq\rho\sqrt{1-e^{-2c/a}},~{}\mathop{\lim\sup}\limits_{t\to\infty}|\tilde{d}(t)|\leq\sqrt{2\eta c/a}.$
[…]
Remark 4.
Based on the facts that $\|\mathbf{R}^{-1}(\theta)\|=1/\ell$ ($\ell<1$), $\|\bm{\tau}_{d}\|\leq\|\bm{\kappa}_{0}(\mathbf{x}_{d})\|\leq\alpha$ (noting (13), (14), and $\|\bm{\Pi}(\mathbf{x}_{d})\|=1$), and $\|\bm{\varpi}\|\leq\hat{d}\leq d_{m}+\sigma$ (noting (34)), it can be easily verified that the adaptive controller (29) is bounded and its bound can be explicitly expressed as
$\|\mathbf{u}\|\leq\dfrac{1}{\ell}(k\rho+\alpha+d_{m}+\delta).$ (45)
This implies that the input constraint $\|\mathbf{u}\|\leq u_{m}$ can be met by judiciously choosing the design parameters such that $k\rho+\alpha+d_{m}+\delta\leq\ell u_{m}$. However, it is noted that achieving smaller convergence bounds for $\mathbf{x}_{e}$ and $\tilde{d}$ requires a larger value of $k$. Thus, a compromise should be reached according to the robot actuation capability and performance requirement.
Remark 5.
As defined in (1), the kinematics of $\theta$ is
$\dot{\theta}=\omega+u_{d2},$ (46)
[…]
IV Simulation Results
In this section, we demonstrate the proposed control method in a $6.4\,\text{m}\times 3.4\,\text{m}$ rectangular workspace $\mathcal{W}^{\ast}$ cluttered with 8 circular obstacles, denoted as $\mathcal{O}_{i}^{\ast}$, whose centers and radii are listed in Table I. The distance between the virtual and actual control points is $\ell=0.05\,\rm m$, the radius of the circumscribed circle around $\bm{P}$ is $r=0.2\,\rm m$, and the clearance constant in Assumption 2 is $h=0.2\,\rm m$. The size of the safety margin is set to $\epsilon=0.1\,\rm m$, and the obstacle influence region is specified by $\epsilon^{\ast}=0.2\,\rm m$. The goal location is taken as $\mathbf{x}^{\ast}=[2.5,1]^{\top}\,\rm m$. Assume that the control input constraint is $u_{m}=1.5$, and the disturbance $\mathbf{u}_{d}$ is of the following form
$\mathbf{u}_{d}=0.01\times\left[\begin{matrix}\sin(0.2t)+1\\ \cos(0.3t)-2\end{matrix}\right].$ (47)
According to Remark 4, the planner parameters are chosen as $\alpha=0.03$ and $\beta=0.005$, while the controller parameters are chosen as $\rho=0.06$, $k=0.1$, $\psi=0.005$, $\eta=0.1$, $\gamma=0.01$, $d_{m}=0.03$, $\delta=0.005$, and $\hat{d}(0)=0.01$. The simulation duration is $500\,\rm s$, and the sample step is $0.01\,\rm s$. We begin by verifying the efficiency of the proposed tangent cone-based path planning method (denoted as “TC-based planner”) in the absence of disturbances. For comparison purposes, the PF-based path planning method presented in [30] (denoted as “PF-based planner”) is also simulated. A slight modification is made for [30] to facilitate its implementation. The attractive and repulsive potentials are chosen as
$U_{att}(\mathbf{x}_{d})=$ $\dfrac{1}{2}k_{a}\|\mathbf{x}_{d}-\mathbf{x}^{\ast}\|^{2},$
$U_{rep}(\mathbf{x}_{d})=$ $\dfrac{1}{2}k_{r}\left(\sum_{i=0}^{8}\dfrac{1}{\rho_{i}}\right)\|\mathbf{x}_{d}-\mathbf{x}^{\ast}\|^{2},$
where $k_{a}$ and $k_{r}$ are positive weights, $\rho_{0}:=1-(x_{d}/2.9)^{20}-(y_{d}/1.4)^{20}$, while $\rho_{i}:=\|\mathbf{x}_{d}-\mathbf{c}_{i}\|^{2}-(r+r_{i}+\epsilon)^{2}$, $i\in\mathbb{I}$. The total velocity is then given by
$\bm{\tau}_{d}=-\nabla U_{att}(\mathbf{x}_{d})-\nabla U_{rep}(\mathbf{x}_{d}).$ (48)
For fair comparison, we chose the parameters of the PF-based path planner to be $k_{a}=0.05$ and $k_{r}=0.0001$, such that the two planners have nearly identical convergence times. Figure 6 shows the obtained robot navigation trajectories of both path planners starting at a set of initial positions (purple points). As can be seen, both planners can successfully guide the virtual control point $\bm{P}$ from different initial positions to the goal (red point), while avoiding all obstacle regions $\mathcal{O}_{i}^{\epsilon}$, $i=1,2,...,8$. However, the proposed planner yields shorter trajectories than the PF-based one. The comparison results of the velocity norm $\|\bm{\tau}_{d}\|$ are depicted in Fig. 7, from which it is clear that the PF-based planner requires larger velocity than the proposed one. Actually, $\bm{\tau}_{d}$ under the proposed planner is upper bounded by $\alpha=0.03$, which is consistent with our analysis in Remark 3. The above results demonstrate the effectiveness and superiority of the proposed path planner.
TABLE I: Geometrical details of the obstacles.
Index Configuration Index Configuration
Center (m) Radius (m) Center (m) Radius (m)
$\mathcal{O}_{1}^{\ast}$ $[-2,-0.55]^{\top}$ $0.10$ $\mathcal{O}_{5}^{\ast}$ $[0.4,0.55]^{\top}$ $0.25$
$\mathcal{O}_{2}^{\ast}$ $[-0.9,0.85]^{\top}$ $0.10$ $\mathcal{O}_{6}^{\ast}$ $[0.7,-0.6]^{\top}$ $0.10$
$\mathcal{O}_{3}^{\ast}$ $[-0.7,-0.5]^{\top}$ $0.35$ $\mathcal{O}_{7}^{\ast}$ $[2,-0.6]^{\top}$ $0.25$
$\mathcal{O}_{4}^{\ast}$ $[-2.1,0.6]^{\top}$ $0.15$ $\mathcal{O}_{8}^{\ast}$ $[1.8,0.7]^{\top}$ $0.15$
Figure 6: Navigation trajectories of the two path planners.
Figure 7: Velocity norms of the two path planners.
We next illustrate the performance of the proposed adaptive tube-following controller (29) (denoted as “TF controller”) in the presence of disturbance given by (47). As a case study, the trajectory starting from the 3nd initial position is selected as the reference trajectory (see Fig. 6). For comparison purposes, the extensively used proportional-integral (PI) controller of the following form is simulated:
$\mathbf{u}=-k_{p}\mathbf{x}_{e}-k_{i}\int_{0}^{t}\mathbf{x}_{e}(\tau)\text{d}\tau,$
where $k_{p}=0.5$ and $k_{i}=0.3$ are constant gains. Furthermore, in order to verify the robustness of the PF-based planner (48) (denoted here as “PF controller”), we execute it in the presence of disturbances and define the difference between the ideal and actual trajectories as the position tracking error $\mathbf{x}_{e}$. The closed-loop responses of these three controllers are shown in Fig. 8. In the top left subfigure of Fig. 8, the tracking error $\mathbf{x}_{e}$ under both the TF and PI controllers is shown to remain within a small neighborhood of the origin (but quantitatively, the TF controller has higher tracking accuracy than the PI controller), while satisfying the tube constraint (24). However, the PF controller yields a larger tracking error and violates the tube constraint. The latter may result in collision with obstacles or the workspace boundary. The WMR’s heading angle is plotted in the top right subfigure of Fig. 8, from which it is clear that $\theta$ under the TF and PI controllers is bounded and converges to a nearly constant value (precisely, the TF controller has a higher stability), while under the PF controller $\theta$ continuously increases. This implies that the PF controller (i.e., path planner (48)) is susceptible to disturbances. The control input norm $\|\mathbf{u}\|$ is shown in the bottom left subfigure of Fig. 8. Obviously, the control input $\mathbf{u}$ of the TF and PI controllers satisfies the magnitude limits $\|\mathbf{u}\|\leq u_{m}$, while the PF controller fails. The parameter estimate $\hat{d}$ is given in the bottom right subfigure of Fig. 8. As can be seen, $\hat{d}$ remains within the set $\Omega_{\delta}$, indicating the effectiveness of the projection adaptive law (31). Finally, the position tracking trajectory of the proposed TF controller (29) is shown in Fig. 9. Intuitively, the virtual control point $\bm{P}$ accurately tracks the reference trajectory, while remaining with the predefined tube. Therefore, the WMR reaches the goal position without colliding with any obstacles (dark gray balls), as clearly seen in Fig. 9. Overall, the proposed integrated path planning and tube-following control scheme can successfully drive the WMR to reach the predefined goal without colliding with any obstacles, despite the presence of input disturbances.
Figure 8: Closed-loop control responses.
Figure 9: Position tracking trajectory with WMR snapshots.
V Experiments
In this section, several experiments are carried out to verify the effectiveness of the proposed method.
V-A Experimental setup
In the experiments, several simple scenarios are considered where the WMR is tasked with navigating towards a designated goal location, while ensuring avoidance of any obstacles en route. The experimental platform is shown in Fig. 10, which consists of a Mona robot [31] with nonholonomic dynamics, several hot obstacles containing iron powders that can heat up themselves, a RGB camera to observe the pose and location of the WMR, a thermal camera to obtain the size and location of the hot obstacles, and a control PC to collect feedback data and send motion commands. The Wi-Fi communication (at a rate of $10\,\rm Hz$) between the control PC and the robot is built by ROS. All experiments are performed in an arena with a size of $\rm 1.8\,m\times 0.8\,m$. It should be emphasized that since the WMR in our lab has a very limited sensing capability, the distance between the WMR and obstacle boundary is obtained in real-time from the RGB and thermal cameras, rather than sensors onboard the WMR. Regarding the robot configuration, we set the radius of the circle surrounding the robot to $r=0.06\,\text{m}$, the parameter for the change of coordinates to $\ell=-0.02$, the safety margin to $\epsilon=0.04\,\text{m}$, and the obstacle influence region to $\epsilon^{*}=0.05\,\text{m}$. In all experiments, the parameters of the path planner are set to $\alpha=0.03$ and $\beta=0.005$, while the controller parameters are set to $\rho=0.03$, $k=1.5$, $\psi=0.005$, $\eta=0.1$, $\gamma=0.01$, $d_{m}=0.036$, $\delta=0.005$, and $d(0)=0.01$.
V-B Experimental results
We here report the results of two representative experiments, where the extracted thermal image and the RGB image are merged to record the experimental process. In the first experiment, six obstacles are placed in the arena. The WMR starts from the left side of the arena (coordinates: $[0.27,0.21]^{\top}\text{m}$), and is tasked with reaching the goal location situated on the right side (coordinates: $[1.73,0.54]^{\top}\text{m}$). The results are shown in Fig. 11, from which it can be observed that the proposed path planner generates a collision-free navigation path for the WMR (i.e., the red line in Fig. 11(a)), and the tube-following controller guarantees that the WMR tracks the desired path along with the predefined safe tube. The latter can be seen from Fig. 11(c), where the norm of the position tracking error is smaller than the tube width $\rho=0.03\text{m}$ all the time. In the second experiment, we add two more hot obstacles to the arena, which makes the environment more complex for the WMR. In this case, the WMR starts from the right side of the arena (coordinate: $[1.83,0.68]^{\top}\text{m}$) and is required to reach a goal location at the left side (coordinate: $[0.20,0.40]^{\top}\text{m}$). The experiment results are shown in Fig. 12, where we can observe similar outcomes as in the first experiment. Our path planner successfully generates a collision-free navigation path for the WMR and the tube-following controller drives the WMR to track the generated path without running out the predefined tube (see Fig. 12(c)). Furthermore, we perform multiple trials with different initial and goal locations to further test the proposed method. From these experiments, we consistently observe the effectiveness of our proposed robot navigation algorithm. More details can be found in the supplementary video (https://vimeo.com/895801720).
Figure 10: The experiment setup: (a) experimental platform; (b) cameras; (c) Mona robot.
Figure 11: The experimental scenario with six obstacles: (a) the generated path (red line) and the actual trajectory (black line) of the WMR; (b) screenshots of the experiment; (c) position error; (d) WMR’s orientation; (e) parameter estimation $\hat{d}$.
Figure 12: The experimental scenario with eight obstacles: (a) the generated path (red line) and the actual trajectory (black line) of the WMR; (b) screenshots of the experiment; (c) position error; (d) WMR’s orientation; (e) parameter estimation $\hat{d}$.
VI Conclusion
In this paper, we have proposed an integrated path planning and tube-following control scheme for collision-free navigation of WMRs in a convex workspace cluttered with multiple spherical obstacles. A path planner was designed using Bouligand’s tangent cones to generate safe, collision-free motion paths for the WMR. The resulting vector field asymptotically guides the WMR from almost all initial configurations in the free space towards the goal location, without colliding with any obstacles along the way. Then, an adaptive tube-following controller was derived, which guarantees that the actual trajectory of the WMR remains within a predefined safe tube around the reference trajectory, despite the presence of disturbances. The tube-following control significantly enhances the safety of robot navigation in the perturbed scenarios. In particular, the proposed closed-loop navigation strategy can be implemented using only the local sensing information. Finally, numerical simulations and experiments illustrate the effectiveness of the proposed autonomous navigation algorithm. Future work will focus on reactive collision-free navigation of multiple WMRs. [Proof of Theorem 2] From (18) and (19), it follows that $\bm{\tau}(\mathbf{x})\in\mathbf{T}_{\mathcal{X}_{\epsilon}}(\mathbf{x})$ for all $\mathbf{x}\in\mathcal{X}_{\epsilon}$, which according to Theorem 1 suggests that the free space $\mathcal{X}_{\epsilon}$ is forward invariant. As previously stated, $\bm{\tau}(\mathbf{x})$ is piecewise continuously differentiable, indicating that it is locally Lipschitz on its domain $\mathcal{X}$ [32]. Moreover, the free space $\mathcal{X}{\epsilon}$, as a compact subset of $\mathcal{X}$, is forward invariant. Then, it follows from [33, Theorem 3.3] that the closed-loop kinematics (11) with $\mathbf{x}(0)\in\mathcal{X}_{\epsilon}$ has a unique solution that is defined for all $t\geq 0$. Subsequently, we show that the unique solution converges to a set of stationary points. To this end, we consider the following LFC:
$W(\mathbf{x})=\dfrac{1}{2}\|\mathbf{x}-\mathbf{x}^{\ast}\|^{2}.$ (49)
For ease of analysis, (18) is rewritten as
$\bm{\tau}(\mathbf{x})=\underbrace{(c(\mathbf{x})\mathbf{I}_{2}+(1-c(\mathbf{x}))\mathbf{\Pi}(\mathbf{x}))}_{\mathbf{\Theta}(\mathbf{x})}\bm{\kappa}_{0}(\mathbf{x}),$ (50)
with
$c(\mathbf{x}):=\left\{\begin{split}&1,&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})>\epsilon^{\ast}~{}\text{or}\\ &&&\text{d}_{\mathcal{O}}(\mathbf{x})\leq\epsilon^{\ast}~{}\text{and}~{}\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})\leq 0,\\ &0,&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})\leq\epsilon^{\ast}~{}\text{and}~{}\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})>0.\end{split}\right.$ (51)
Taking the time derivative of $W(\mathbf{x})$ along (11) and noting (13) and (50), one gets
$\dot{W}(\mathbf{x})=-k_{0}(\mathbf{x})(\mathbf{x}-\mathbf{x}^{\ast})^{\top}\mathbf{\Theta}(\mathbf{x})(\mathbf{x}-\mathbf{x}^{\ast}).$ (52)
Since $k_{0}(\mathbf{x})>0$ and $\mathbf{\Pi}(\mathbf{x})$ is positive semidefinite, one has
$\dot{W}(\mathbf{x})\leq 0,~{}\forall\mathbf{x}\in\mathcal{X}_{\epsilon},$ (53)
indicating that $\mathbf{x}=\mathbf{x}^{\ast}$ is a stable equilibrium of (11). From LaSalle’s invariance principle, it follows that the solution of (11) asymptotically converges to the set $\{\mathbf{x}\in\mathcal{X}_{\epsilon}\mid\dot{W}(\mathbf{x})=0\}$, i.e., a set of stationary points. One can infer from (19) and (52) that the stationary points correspond to either $\mathbf{x}=\mathbf{x}^{\ast}$ or the points satisfying $\text{d}_{\mathcal{O}}(\mathbf{x})=\epsilon$ and $\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})/\|\bm{\kappa}_{0}(\mathbf{x})\|=1$. Therefore, for any $\mathbf{x}(0)\in\mathcal{X}_{\epsilon}$, the solution of (11) converges to the set $\{\mathbf{x}^{\ast}\}\cup_{i=1}^{n}\mathcal{S}_{i}$. The set contains $n+1$ stationary points, all of which are isolated due to Assumption 2. Geometrically, the undesired stationary point $\mathbf{s}_{i}\in\mathcal{S}_{i}$ (on the boundary of $\mathcal{O}_{i}^{\epsilon}$) and $\mathbf{x}^{\ast}$ are collinear with $\mathbf{c}_{i}$, but located on the opposite sides of $\mathbf{c}_{i}$, as seen in Fig. 13. As per (52), the analytical form of $\mathbf{s}_{i}$ is obtained as follows:
$\mathbf{s}_{i}=(1+\alpha_{i})\mathbf{c}_{i}-\alpha_{i}\mathbf{x}^{\ast},$ (54)
with $\alpha_{i}:=(r+r_{i}+\epsilon)\|\mathbf{x}^{\ast}-\mathbf{c}_{i}\|^{-1}$.
Figure 13: Illustration of the undesired stationary point $\mathbf{s}_{i}$.
The isolated stationary points $\mathbf{s}_{i}$, $i\in\mathbb{I}$ may prevent us from achieving the objective (motion-to-goal). To check the stability of $\mathbf{s}_{i}$, we define a small neighborhood of $\mathbf{s}_{i}$, i.e., an open ball $\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$ with $0<\epsilon^{\prime}\leq\min\{\epsilon,\epsilon^{\ast}-\epsilon\}$, such that $\bf{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})>0$ holds for all $\mathbf{x}\in\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$. When restricted on $\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$, the feasible set of initial configurations is $\mathcal{F}_{i}:=\mathcal{X}_{\epsilon}\cap\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$, and $\mathbf{x}$ evolves according to
$\dot{\mathbf{x}}=-k_{0}(\mathbf{x})(\mathbf{I}_{2}-\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))\mathbf{b}(\mathbf{x})\mathbf{b}^{\top}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast}).$ (55)
Further define two sets
$\mathcal{M}_{i}:=$$~{}\left\{\mathbf{x}\in\mathcal{X}_{\epsilon}\bigg{|}\text{d}_{\mathcal{O}_{i}}(\mathbf{x})\geq\epsilon,~{}\dfrac{\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})}{\|\bm{\kappa}_{0}(\mathbf{x})\|}=1\right\},$ (56)
$\mathcal{N}_{i}:=$ $~{}\partial\mathcal{X}_{\epsilon}\cap\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime}).$ (57)
Geometrically speaking, all elements of $\mathcal{M}_{i}$ form a radial line extending outward from the stationary point $\mathbf{s}_{i}$ along $\mathbf{s}_{i}-\mathbf{c}_{i}$ (see the yellow line in Fig. 13), while all elements of $\mathcal{N}_{i}$ form a curve, which is a portion of the boundary $\partial\mathcal{O}_{i}^{\epsilon}$ (see the blue curve in Fig. 13). In the following, three cases are considered: 1) $\mathbf{x}(0)\in\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})\cap\mathcal{M}_{i}$; 2) $\mathbf{x}(0)\in\mathcal{N}_{i}\setminus\{\mathbf{s}_{i}\}$; and 3) $\mathbf{x}(0)\in\mathcal{F}_{i}\setminus(\mathcal{M}_{i}\cup\mathcal{N}_{i})$. Case 1): $\mathbf{x}(0)\in\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})\cap\mathcal{M}_{i}$. Note that for all $\mathbf{x}\in\mathcal{M}_{i}$, $\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})/\|\bm{\kappa}_{0}(\mathbf{x})\|=1$ holds, whereby (55) reduces to
$\dot{\mathbf{x}}=-k_{0}(\mathbf{x})(1-\phi(\text{d}_{\mathcal{O}}(\mathbf{x})))(\mathbf{x}-\mathbf{x}^{\ast}).$ (58)
From (20) and (56), it is evident that $0\leq\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))\leq 1$ on the set $\mathcal{M}_{i}$ and $\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))=1$ only when $\mathbf{x}=\mathbf{s}_{i}$. Moreover, $k_{0}(\mathbf{x})$ is strictly larger than 00 by construction. Hence, for all $\mathbf{x}\in\mathcal{M}_{i}\setminus\{\mathbf{s}_{i}\}$, the control vector $\bm{\tau}(\mathbf{x})$ (i.e., the right-hand side of (58)) points directly towards $\mathbf{s}_{i}$ and becomes zero when $\mathbf{x}=\mathbf{s}_{i}$, indicating that the set $\mathcal{M}_{i}$ is forward invariant. Consider the following LFC:
$W_{i}(\mathbf{x})=\dfrac{1}{2}\|\mathbf{x}-\mathbf{s}_{i}\|^{2}.$ (59)
The time derivative of $W_{i}(\mathbf{x})$ along (58) is given by
$\dot{W}_{i}(\mathbf{x})=-k_{0}(\mathbf{x})(1-\phi(\text{d}_{\mathcal{O}}(\mathbf{x})))(\mathbf{x}-\mathbf{s}_{i})^{\top}(\mathbf{x}-\mathbf{x}^{\ast}).$ (60)
Since $(\mathbf{x}-\mathbf{s}_{i})^{\top}(\mathbf{x}-\mathbf{x}^{\ast})\geq 0$ for all $\mathbf{x}\in\mathcal{M}_{i}$, one can conclude from (60) that $\dot{W}_{i}(\mathbf{x})\leq 0$ on the set $\mathcal{M}_{i}$ and $\dot{W}_{i}(\mathbf{x})=0$ only occurs at $\mathbf{x}=\mathbf{s}_{i}$. This implies that $\mathcal{M}_{i}$ is a stable manifold of $\mathbf{s}_{i}$ and any initial state on it will converge to $\mathbf{s}_{i}$. Thus, all $\mathbf{x}(0)\in\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})\cap\mathcal{M}_{i}\subset\mathcal{M}_{i}$ will converge to $\mathbf{s}_{i}$. Case 2): $\mathbf{x}(0)\in\mathcal{N}_{i}\setminus\{\mathbf{s}_{i}\}$. For all $\mathbf{x}\in\mathcal{N}_{i}\setminus\{\mathbf{s}_{i}\}$, we have $\text{d}_{\mathcal{O}}(\mathbf{x})=\epsilon$ and $\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))=1$. As such, (55) becomes
$\dot{\mathbf{x}}=-k_{0}(\mathbf{x})\mathbf{Q}(\mathbf{b}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast}),$ (61)
where $\mathbf{Q}(\mathbf{b}(\mathbf{x})):=\mathbf{I}_{2}-\mathbf{b}(\mathbf{x})\mathbf{b}^{\top}(\mathbf{x})$ is defined for notational brevity. Consider again the LFC $W_{i}(\mathbf{x})$ in (59). Then, taking the time derivative of $W_{i}$ along (61) and noting (54) and the fact that $(\mathbf{x}-\mathbf{c}_{i})^{\top}\mathbf{Q}(\mathbf{b}(\mathbf{x}))=0$, we get
$\dot{W}_{i}(\mathbf{x})=$ $~{}k_{0}(\mathbf{x})\alpha_{i}(\mathbf{x}-\mathbf{x}^{\ast})^{\top}\mathbf{Q}(\mathbf{b}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast})$
$-k_{0}(\mathbf{x})(1-\alpha_{i})(\mathbf{x}-\mathbf{c}_{i})^{\top}\mathbf{Q}(\mathbf{b}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast})$
$=$ $~{}k_{0}(\mathbf{x})\alpha_{i}(\mathbf{x}-\mathbf{x}^{\ast})^{\top}\mathbf{Q}(\mathbf{b}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast})$
$=$ $~{}k_{0}(\mathbf{x})\alpha_{i}(1-\cos\psi)\|\mathbf{x}-\mathbf{x}^{\ast}\|^{2}>0,$ (62)
where $\psi$ is the angle between the vectors $\mathbf{b}(\mathbf{x})$ and $\bm{\kappa}_{0}(\mathbf{x})$ and satisfies $20<\psi<\pi/$ 2 for all $\mathbf{x}\in\mathcal{N}_{i}\setminus\{\mathbf{s}_{i}\}$. Further consider the following LFC:
$V_{i}(\mathbf{x})=\frac{1}{2}\|\mathbf{x}-\mathbf{c}_{i}\|^{2},$ (63)
whose time derivative along (61) is given by
$\dot{V}_{i}(\mathbf{x})=-k_{0}(\mathbf{x})(\mathbf{x}-\mathbf{c}_{i})^{\top}\mathbf{Q}(\mathbf{b}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast})=0,$ (64)
showing that for all $\mathbf{x}\in\mathcal{N}_{i}\setminus\{\mathbf{s}_{i}\}$, the control vector is tangent to the boundary of $\mathcal{O}_{i}^{\epsilon}$, thus directing the robot to move along $\partial\mathcal{O}_{i}^{\epsilon}$. In view of (VI) and (64), we can conclude that if $\mathbf{x}(0)\in\mathcal{N}_{i}\setminus{\mathbf{s}_{i}}$, then $\mathbf{x}$ will keep away from $\mathbf{s}_{i}$ along $\partial\mathcal{O}_{i}^{\epsilon}$ until it leaves the set $\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$. Case 3): $\mathbf{x}(0)\in\mathcal{F}_{i}\setminus(\mathcal{M}_{i}\cup\mathcal{N}_{i})$. Actually, for any $\mathbf{x}\in\mathcal{F}_{i}\setminus(\mathcal{M}_{i}\cup\mathcal{N}_{i})$, it holds that $\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))\in(0,1)$, and $\bf{\kappa}_{0}(\mathbf{x})$ and $\mathbf{b}(\mathbf{x})$ are not collinear. Moreover, we can decompose $\bm{\tau}(\mathbf{x})=\mathbf{\Pi}(\mathbf{x})\bm{\kappa}{0}(\mathbf{x})$ into two components: the component $\bm{\tau}(\mathbf{x})_{\parallel}=\mathbf{b}(\mathbf{x})\mathbf{b}^{\top}(\mathbf{x})\bm{\tau}(\mathbf{x})$ parallel to $\mathbf{b}(\mathbf{x})$ and the component $\bm{\tau}(\mathbf{x})_{\perp}=(\mathbf{I}_{2}-\mathbf{b}(\mathbf{x})\mathbf{b}^{\top}(\mathbf{x}))\bm{\tau}(\mathbf{x})$ perpendicular to $\mathbf{b}(\mathbf{x})$, as shown in Fig. 13. With (19) in mind, one can verify that
$\mathbf{b}^{\top}(\mathbf{x})\bm{\tau}(\mathbf{x})_{\parallel}=$ $~{}\mathbf{b}^{\top}(\mathbf{x})\bm{\tau}(\mathbf{x})$
$=$ $~{}(1-\phi(\text{d}_{\mathcal{O}}(\mathbf{x})))\mathbf{b}^{\top}(\mathbf{x})\bm{\kappa}_{0}(\mathbf{x})>0,$ (65)
which implies that $\bm{\tau}(\mathbf{x}){\parallel}$ and the bearing vector $\mathbf{b}(\mathbf{x})$ have the same direction, causing $\bm{\tau}(\mathbf{x}){\parallel}$ to point towards $\mathbf{c}_{i}$. A straightforward calculation yields $(\mathbf{I}_{2}-\mathbf{b}(\mathbf{x})\mathbf{b}^{\top}(\mathbf{x})\mathbf{\Pi}(\mathbf{x})=\mathbf{Q}(\mathbf{b}(\mathbf{x}))$, which allows us to express $\bm{\tau}(\mathbf{x})_{\perp}=\mathbf{Q}(\mathbf{b}(\mathbf{x}))\bm{\kappa}_{0}(\mathbf{x})$. Under $\bm{\tau}(\mathbf{x})_{\perp}$, (55) turns to be (61), which results in $\dot{W}_{i}(\mathbf{x})>0$ as seen in (VI). Thus, $\bm{\tau}(\mathbf{x})_{\perp}$ generates a velocity component tangent to $\mathbf{b}$, causing $\mathbf{x}$ to move away from $\mathbf{s}_{i}$. Since $\bm{\tau}(\mathbf{x})$ and the vector $\mathbf{s}_{i}-\mathbf{x}$ are situated on opposite sides of $\mathbf{b}(\mathbf{x})$ (see Fig. 13), $\mathbf{x}$ cannot move toward $\mathbf{s}_{i}$ on the set $\mathcal{F}_{i}\setminus(\mathcal{M}_{i}\cup\mathcal{N}_{i})$. From Fig. 13, one can intuitively observe that $\bm{\tau}(\mathbf{x})$ guides $\mathbf{x}$ towards the boundary of $\mathcal{O}_{i}^{\epsilon}$ (attributed to $\bm{\tau}(\mathbf{x})_{\parallel})$, while simultaneously keeping away from the manifold $\mathcal{M}_{i}$ (attributed to $\bm{\tau}(\mathbf{x})_{\perp}$). Therefore, for any $\mathbf{x}(0)\in\mathcal{F}_{i}\setminus(\mathcal{M}_{i}\cup\mathcal{N}_{i})$, the solution of (11) will either directly leaves the ball $\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$, or first converges to the set $\mathcal{N}_{i}\setminus\{\mathbf{s}_{i}\}$ and then leaves the ball $\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$ along $\partial\mathcal{O}_{i}^{\epsilon}$ (as shown in Case 2). The three cases above demonstrate that each stationary point $\mathbf{s}_{i}$ ($i\in\mathbb{I}$) is an unstable fixed point, but there exists one line of initial conditions, namely the stable manifold $\mathcal{M}_{i}$, that is attracted to $\mathbf{s}_{i}$. Note that $\mathcal{M}_{i}$ is a 1-dimensional manifold with boundary and thus has zero measure [34]. As such, $\mathbf{x}^{\ast}$ is an almost globally asymptotically stable equilibrium, with a basin of attraction consisting of the free space $\mathcal{X}_{\epsilon}$, except for a set of measure zero. Furthermore, as $\mathbf{x}^{\ast}\in\text{int}(\mathcal{X}_{\epsilon})$, there certainly exists $r^{\ast}>0$ such that $\text{d}_{\mathcal{O}}(\mathbf{x})>\epsilon^{\ast}$ for all $\mathbf{x}\in\mathcal{B}(\mathbf{x}^{\ast},r^{\ast})$. On the ball $\mathcal{B}(\mathbf{x}^{\ast},r^{\ast})$, the kinematics reduces to
$\dot{\mathbf{x}}=-k_{0}(\mathbf{x})(\mathbf{x}-\mathbf{x}^{\ast}),$ (66)
It can be easily verified that $k_{0}(\mathbf{x})\geq\alpha/\sqrt{(r^{\ast})^{2}+\beta^{2}}$ on the ball $\mathcal{B}(\mathbf{x}^{\ast},r^{\ast})$. Given this, (66) implies the local exponential stability of $\mathbf{x}=\mathbf{x}^{\ast}$. Furthermore, by using the inequality $|z|-z^{2}/\sqrt{z^{2}+\beta^{2}}\leq\beta$, $\forall z\in\mathbb{R}$, (52) reduces to
$\dot{W}_{i}(\mathbf{x})\leq$ $-\alpha\|\mathbf{x}-\mathbf{x}^{\ast}\|+\alpha\beta$
$\leq$ $-\sqrt{2}\alpha W_{i}(\mathbf{x})^{\frac{1}{2}}+\alpha\beta,$ (67)
[…]

### v2

@@
-Collision-Free Navigation of Wheeled Mobile Robots: An Integrated Path Planning and Tube-Following Control Approach
+InPTC: Integrated Planning and Tube-Following Control for Prescribed-Time Collision-Free Navigation of Wheeled Mobile Robots
 Abstract
-In this paper, an integrated path planning and tube-following control scheme is proposed for collision-free navigation of a wheeled mobile robot (WMR) in a compact convex workspace cluttered with sufficiently separated spherical obstacles.
-An analytical path planning algorithm is developed based on Bouligand’s tangent cones and Nagumo’s invariance theorem, which enables the WMR to navigate towards a designated goal location from almost all initial positions in the free space, without entering into augmented obstacle regions with safety margins.
-We further construct a virtual “safe tube” around the reference trajectory, ensuring that its radius does not exceed the size of the safety margin.
-Subsequently, a saturated adaptive controller is designed to achieve safe trajectory tracking in the presence of disturbances.
-It is shown that this tube-following controller guarantees that the WMR tracks the reference trajectory within the predefined tube, while achieving uniform ultimate boundedness of both the position tracking and parameter estimation errors.
-This indicates that the WMR will not collide with any obstacles along the way.
-Finally, we report simulation and experimental results to validate the effectiveness of the proposed method.
+In this article, we propose a novel approach, called InPTC (Integrated Planning and Tube-Following Control), for prescribed-time collision-free navigation of wheeled mobile robots in a compact convex workspace cluttered with static, sufficiently separated, and convex obstacles.
+A path planner with prescribed-time convergence is presented based upon Bouligand’s tangent cones and time scale transformation (TST) techniques, yielding a continuous vector field that can guide the robot from almost all initial positions in the free space to the designated goal at a prescribed time, while avoiding entering the obstacle regions augmented with safety margin.
+By leveraging barrier functions and TST, we further derive a tube-following controller to achieve robot trajectory tracking within a prescribed time less than the planner’s settling time.
+This controller ensures the robot moves inside a predefined “safe tube” around the reference trajectory, where the tube radius is set to be less than the safety margin.
+Consequently, the robot will reach the goal location within a prescribed time while avoiding collision with any obstacles along the way.
+The proposed InPTC is implemented on a Mona robot operating in an arena cluttered with obstacles of various shapes.
+Experimental results demonstrate that InPTC not only generates smooth collision-free reference trajectories that converge to the goal location at the preassigned time of $250\,\rm s$ (i.e., the required task completion time), but also achieves tube-following trajectory tracking with tracking accuracy higher than $0.01\rm m$ after the preassigned time of $150\,\rm s$.
+This enables the robot to accomplish the navigation task within the required time of $250\,\rm s$.
 Index Terms:
-Tube-following control, adaptive control, path planning, robot navigation, wheeled mobile robots.
+Prescribed-time control, path planning, trajectory tracking, collision avoidance, wheeled mobile robots.
 I Introduction
-The path planning and control of wheeled mobile robots (WMRs) is an important problem in robotics due to the extensive real-world applications that these systems have, such as cargo transportation, exploration of hazardous environments, automated patrolling [1], to name a few cases.
-In practical applications, WMRs usually operate in obstacle-cluttered environments, which limits the implementation of traditional motion control methods, thus, motivates the development of collision-free navigation algorithms that can steer WMRs from an initial position to a target goal without colliding with any obstacles along the way [2].
-Collision-free robot navigation has garnered significant attention from the robotics and control research communities [3].
-Existing solutions to the collision-free navigation problem of robots can be primarily classified into two categories [4]: global (map-based) methods and local (reactive) methods.
-The former relies on global knowledge of the environment, requiring a priori information about the obstacles (e.g., position and shape).
-In contrast, the latter utilizes only local knowledge of obstacles in the immediate vicinity of the robot, obtained from on-board sensors.
-Among the global methods, a simple and computationally efficient solution is the artificial potential field (APF)-based approach [5], which uses a field of virtual potential forces to push the robot towards a target position and pull it away from obstacles; Representative APF-based methods can be found in [6, 7, 8] and references therein.
-These kind of methods often reach local minima, which hinders achieving global convergence to a designated goal (especially in topologically complex settings).
-Although navigation functions developed in [9] can help to deal with this problem, they come with extra drawbacks, such as the need for unknown tuning of parameters.
-Other global methods for collision-free navigation are e.g., heuristic approaches like $A^{\star}$ [10, 11], rapidly exploring random tree (RRT) [12], genetic algorithms [13], etc.
-As most of them are search-based solutions, their performance can be significantly affected by the scale of the problem [14].
-Optimization-based methods [15], while providing alternatives, typically rely on numerical solutions of the constrained optimization problems that result on heavily computational costs.
-Both the search-based and optimization-based methods are inherently open loop, which makes them vulnerable to disturbances and sensor noise.
-Thus, the actual robot trajectory may deviates from the desired path, hence, risking collisions with the environment.
-Local methods that offer reactive solutions for collision-free robot navigation are highly desirable in practice.
-Particularly in autonomous exploration applications, where WMRs often have limited access to global information and must rely on local sensory data to detect obstacles and perform navigate.
-Bug algorithms that evolved from maze solving algorithms are one of the simplest reaction navigation approaches for mobile robots in planar environments [16, 17].
-Lionis et al.
-[18] proposed locally computable navigation functions for robotic navigation in unknown sphere worlds.
-Filippidis and Kyriakopoulos [19] introduced adjustable navigation functions that are capable of autonomously tuning parameters when new obstacles are discovered.
-Paternain and Ribeiro [20] provided a stochastic extension of these navigation functions with provable convergence properties.
-However, one caveat is that these reactive function-based methods also suffer from local minima [4].
-Arslan and Koditschek [21] proposed a sensor-based reactive method based on a robot-centric spatial decomposition for collision-free navigation with convex obstacles.
+The motion control of wheeled mobile robots (WMRs) is a benchmark problem in robotics due to its key role in extensive real-world applications, such as cargo transportation, automated patrolling, and exploration of hazardous environments [1].
+In these applications, WMRs typically operate in obstacle-cluttered environments, which motivates the development of safe navigation algorithms that can steer WMRs from an initial position to a desired goal location without colliding with any obstacles along the way [2].
+Collision-free navigation of WMRs have garnered significant attention from the robotics and control research communities [3].
+Existing solutions to the robot navigation problem can be primarily classified into two categories [4]: global (map-based) methods and local (reactive) methods.
+The former requires a priori information of environments (e.g., position and shape of obstacles), while the latter utilizes only local knowledge of obstacles obtained from on-board sensors.
+Among the global methods, a computationally efficient solution is the artificial potential field (APF)-based approach [5], which uses a field of potential forces to push the robot towards a goal position and pull it away from obstacles [6, 7, 8].
+Control barrier function (CBF) combined with quadratic programming (CBF-QP) also provides an effective strategy for developing controllers that can both stabilize a system and ensure safety, making it particularly useful for autonomous vehicles operating in obstacle-cluttered environments [9, 10, 11].
+Note that both the APF- and CBF-based methods often reach local minima, which hinders achieving global convergence to the designated goal, especially in topologically complex settings.
+Although navigation functions presented in [12] can overcome this issue, they require unknown tuning of parameters.
+Other global methods include heuristic approaches such as $A^{\star}$ [13], rapidly exploring random tree (RRT) [14], and genetic algorithms [15].
+Since these methods are search-based solutions, their performance is directly affected by the problem scale [16].
+Optimization-based methods [17] provide alternatives, but typically require numerically solving constrained optimization problems, which results in high computational costs.
+Local methods that offer reactive solutions for collision-free robot motions are highly desirable, particularly in autonomous exploration applications, where robots often have limited access to environment information and must rely on local sensory data to detect obstacles.
+Bug algorithm is a simple reaction planning approaches for mobile robots [18, 19].
+Arslan and Koditschek [20] proposed a reactive method based on a robot-centric spatial decomposition for collision-free navigation.
 Huber et al.
-[22] adopted contraction-based dynamical systems theory to achieve a closed-form solution to the avoidance of moving obstacles.
-Recently, reinforcement learning methods were used in [23], however, the learning process requires a substantial amount of interaction with the real world.
-Note that although most of the aforementioned reactive approaches are closed-form, they only consider nominal (holonomic or non-holonomic) robot kinematics, without taking into account disturbances/uncertainties that may arise from external forces, modeling errors, inaccurate controls, sensor noise, data dropout, etc.
-These unknown actions/terms inevitably degrade the system’s performance, potentially leading to deviations of the desired trajectory, thus, jeopardizing the robot’s safety.
-In this paper, we study the collision-free navigation problem of differential-drive WMRs operating in a compact convex workspace cluttered with multiple sufficiently separated spherical obstacles.
-To overcome the non-holonomic constraints inherent in this type of systems, an off-center point is selected as the virtual control point, enabling to transform the original non-holonomic kinematics into a fully-actuated two degrees-of-freedom system [24, 25].
-Building upon this model, we propose a novel robot navigation scheme that integrates a path planner and a tube-following controller.
-By “tube-following” it is meant that the tracking trajectory of the WMR should remain within a virtual safe tube around the reference trajectory to ensure its motion safety.
-The proposed method can achieve collision-free navigation in cluttered environments from almost all initial configurations in the robot’s free space despite the presence of disturbances.
-To the best of the authors’ knowledge, this is the first time an integrated path planning and tube-following control scheme is proposed for safe robot navigation in cluttered environments.
-The effectiveness of the proposed method is validated through numerical simulations and experiments.
-Specifically, the original contributions of the paper are two-fold:
+[21] adopted contraction-based dynamical systems theory to achieve dynamic obstacle avoidance.
+Berkane et al.
+[4] proposed a sensor-based navigation algorithm that employs Nagumo’s invariance theorem to guarantee robot safety by projecting the nominal velocity onto safe velocity cones (e.g., Bouligand’s tangent cones) when the robot is close to obstacle boundaries.
+Although most reactive approaches offer closed-form solutions, they typically consider only the nominal robot kinematics, without accounting for disturbances that may arise, for example, from sensor noise, imperfect wheel alignment, and motor or actuator asymmetry.
+These unmodeled factors inevitably degrade system performance, potentially leading to deviations of the desired trajectory and compromising the robot’s safety.
+Furthermore, the aforementioned methods can only guarantee that WMRs reach the desired goal asymptotically, which is undesirable for time-critical missions.
+There has been extensive research on the trajectory tracking control of WMRs.
+Zhai and Song [22] combined the adaptive and sliding mode control techniques to address the trajectory tracking problem of WMRs.
+Zheng et al.
+[23] presented an adaptive sliding mode control method for trajectory tracking of WMRs, and introduced a barrier function-like control gain to prevent input saturation.
+Kim and Singh[24] proposed a differential-flatness-based feedback controller to achieve the tracking of prescribed point-to-point trajectories.
+Recently, Zhou et al.[25] proposed a homogeneous control strategy to solve the trajectory tracking problem for perturbed unicycle mobile robots, where the convergence rate can be tuned by selecting a proper homogeneity degree.
+While these methods are effective in addressing stability and steady-state performance, they often do not explicitly consider the critical requirement of maintaining the WMR’s actual motion within a predefined safe tube around the reference trajectory.
+This actually imposes constraints on the tracking errors.
+Furthermore, most of these methods can only achieve asymptotic position tracking, and they typically lack the capability to preassign the convergence time according to the task requirements.
+This paper examines the planning and control problems for prescribed-time collision-free navigation of WMRs in a convex workspace cluttered with static, sufficiently separated, and convex obstacles.
+By selecting an off-axis point as the virtual control point, the non-holonomic kinematics of WMRs is transformed into a fully-actuated system [26, Section 11.6.1]).
+Building upon this model, we propose a novel approach, called InPTC (Integrated Planning and Tube-Following Control), to achieve collision-free navigation of WMRs within a finite time that can be preassigned according to the task requirement, even in the presence of disturbances.
+Numerical simulations and experiments illustrate the effectiveness of the proposed InPTC.
+The main contributions of this paper are twofold:
   1. 1)
-    Inspired by [4], a safe path planner is presented based on Bouligand’s tangent cones [26] and Nagumo’s invariance theorem [27], yielding a continuous saturated vector field that can guide the WMR from almost all initial positions in the free space towards a designated goal, without entering into the augmented obstacle regions with safety margins.
-In essence, the proposed planning algorithm is sensor-based, and the vector field can be generated in real-time using only the local sensing information (i.e., the distance between the WMR and the obstacle boundaries).
+    A reactive path planner with prescribed-time convergence is developed based on Bouligand’s tangent cones [4] and time scale transformation [27].
+This planner yields a continuous vector field that can guide the WMR from almost all initial positions in the free space to reach the goal at a prescribed finite time, while avoiding entering the obstacle regions augmented by safety margins.
   2. 2)
-    We propose an adaptive tube-following control method that defines a virtual safe tube around the reference trajectory with its radius no larger than the size of the safety margin.
-It is shown that the derived controller ensures the tracking of the reference trajectory within the predefined safe tube, while achieving uniform ultimate boundedness of both the trajectory tracking and parameter estimation errors, despite the presence of disturbances.
-By design, the chosen tube radius ensures that the WMR’s trajectory never enters into the actual obstacle regions, thus achieving the collision-free robot navigation in the perturbed conditions.
+    By leveraging time scale transformation and barrier functions, a prescribed-time tube-following controller is derived for reference trajectory tracking.
+This controller guarantees that the tracking error converges to a small residual set within a prescribed finite time, whilst the WMR’s position remains within a predefined safe tube around the reference trajectory, despite the presence of disturbances.
+The remainder of the paper is organized as follows: In Section II, we describe the kinematics and operating environment of WMRs, and formulate the prescribed-time safe navigation problem.
+In Section III, an integrated planning and control framework is proposed to address this problem.
+The simulation and experimental results are presented in Sections IV and V, respectively.
+Finally, Section VI provides concluding remarks and discusses future work.
 […]
@@
 $\left[\begin{matrix}\dot{\bar{x}}\\ \dot{\bar{y}}\\ \dot{\bar{\theta}}\end{matrix}\right]=\left[\begin{matrix}\cos\bar{\theta}&0\\ \sin\bar{\theta}&0\\ 0&1\end{matrix}\right](\mathbf{u}+\mathbf{u}_{d}),$ (1)
-where $\mathbf{u}:=[v,\omega]^{\top}\in\mathbb{R}^{2}$ is the the control input vector with $v$ and $\omega$ being the linear and angular velocities of the WMR, respectively, and $\mathbf{u}_{d}\in\mathbb{R}^{2}$ is the matched disturbance that may arise from control errors, sensor noise, etc.
+where $\mathbf{u}:=[v,\omega]^{\top}\in\mathbb{R}^{2}$ is the the control input vector with $v$ and $\omega$ being the linear and angular velocities of the WMR, respectively, and $\mathbf{u}_{d}\in\mathbb{R}^{2}$ is the matched disturbance that may arise from sensor noise, imperfect wheel alignment, and motor or actuator asymmetry.
 Figure 1: Sketch of the wheeled mobile robot.
-To facilitate the subsequent design and analysis, we select an off-center point $\bm{P}$ as the virtual control point, which locates at a distance $\ell\neq 0$ away from $\bar{\bm{P}}$ along the longitudinal axis of the WMR [24, 25], as shown in Fig.
+To facilitate the subsequent design and analysis, we select an off-axis point $\bm{P}$ as the virtual control point, which locates at a distance $\ell\neq 0$ away from $\bar{\bm{P}}$ along the longitudinal axis of the WMR [26], as shown in Fig.
 1.
@@
 $\dot{\mathbf{x}}=\mathbf{R}(\theta)(\mathbf{u}+\mathbf{u}_{d}),$ (3)
-where $\mathbf{R}(\theta)\in\mathbb{R}^{2\times 2}$ is defined as
-$\mathbf{R}(\theta):=\left[\begin{matrix}\cos\theta&-\ell\sin\theta\\ \sin\theta&\ell\cos\theta\\ \end{matrix}\right],$ (4)
-which is full rank. By setting $|\ell|\leq 1$, it follows that $\|\mathbf{R}(\theta)\|=\max\{1,|\ell|\}=1$ and $\|\mathbf{R}^{-1}(\theta)\|=\max\{1,1/|\ell|\}=1/\ell$.
+where $\mathbf{R}(\theta):=[\cos\theta,-\ell\sin\theta;\sin\theta,\ell\cos\theta]$ is full rank.
+This enables us to freely steer the WMR’s position regardless of nonholonomic constraints.
+By setting $|\ell|\leq 1$, it follows that $\|\mathbf{R}(\theta)\|=\max\{1,|\ell|\}=1$.
 Assumption 1.
@@
 Remark 1.
-The full rank of $\mathbf{R}(\theta)$ enables us to freely control the position of the WMR (specifically, the virtual control point $\bm{P}$) without being restricted by nonholonomic constraints.
+The kinematics given by (1) represents a general kinematic model for unicycle mobile robots, such as differential drive robots and synchro drive robots.
+Additionally, following the input/output linearization outlined in [26, Section 11.6.1], the nonholonomic kinematics (1) is transformed into a full-actuated system (3), via a change of coordinates given by (2).
+This transformed model can be rewritten in the form of a single integrator system, which has the same structure as the kinematics of omnidirectional robots with holonomic constraints.
+Consequently, the proposed method building upon (3) is applicable for a broad class of WMRs.
 II-B Operating environment
-Consider a WMR operating inside a closed compact convex workspace $\mathcal{W}^{\ast}\subset\mathbb{R}^{2}$, punctured by a set of $n\in\mathbb{N}$ obstacles $\mathcal{O}_{i}^{\ast}$, $i\in\mathbb{I}:=\{1,2,...,n\}$, which are represented by open balls with centers $\mathbf{c}_{i}\in\mathbb{R}^{2}$ and radii $r_{i}>0$.
+Consider a WMR operating inside a closed compact convex workspace $\mathcal{W}^{\ast}\subset\mathbb{R}^{2}$, punctured by a set of $n$ static obstacles $\mathcal{O}_{i}^{\ast}$, $i\in\mathbb{I}:=\{1,2,...,n\}$, which are represented by open balls with centers $\mathbf{c}_{i}\in\mathbb{R}^{2}$ and radii $r_{i}>0$.
 A circumscribed circle centered at $\bm{P}$ with radius $r>0$ is constructed, which is the smallest circle enclosing the WMR (see Fig.
 1).
-To ensure that the WMR can navigate freely between any of the obstacles in $\mathcal{W}^{\ast}$, we make the following assumption:
+To ensure that the WMR can navigate freely between any of the obstacles in $\mathcal{W}^{\ast}$, we make the following assumption [4, 20]:
 Assumption 2.
@@
 where $h>0$ is a constant.
-For ease of design, the WMR is considered as a point (i.e., the virtual control point $\bm{P}$), by transferring the volume of the circumscribed circle to the other workspace entities.
-Then, for the point $\bm{P}$, the workspace is defined as
-$\mathcal{W}:=\{\mathbf{x}\in\mathbb{R}^{2}\mid\text{d}_{\complement\mathcal{W}^{\ast}}(\mathbf{x})\geq r\},$ (7)
-and the obstacle regions are defined as the following pairwise disjoint spherical subsets of $\mathbb{R}^{2}$:
-$\mathcal{O}_{i}:=\{\mathbf{x}\in\mathbb{R}^{2}\mid\beta_{i}(\mathbf{x})<0\},~{}i\in\mathbb{I},$ (8)
-where $\beta_{i}(\mathbf{x}):=\|\mathbf{x}-\mathbf{c}_{i}\|-(r+r_{i})$.
+For ease of design, the WMR is considered as a point (i.e., the off-axis point $\bm{P}$), by transferring the volume of the circumscribed circle to the other workspace entities.
+For the point $\bm{P}$, the workspace is $\mathcal{W}:=\{\mathbf{x}\in\mathbb{R}^{2}\mid\text{d}_{\complement\mathcal{W}^{\ast}}(\mathbf{x})\geq r\}$, and the obstacle regions are $\mathbb{R}^{2}$: $\mathcal{O}_{i}:=\{\mathbf{x}\in\mathbb{R}^{2}\mid\beta_{i}(\mathbf{x})<0\},~{}i\in\mathbb{I}$, where $\beta_{i}(\mathbf{x}):=\|\mathbf{x}-\mathbf{c}_{i}\|-(r+r_{i})$.
 Thus, the free space of $\bm{P}$ is given by the closed set $\mathcal{X}:=\mathcal{W}\setminus\mathcal{O}$ with $\mathcal{O}:=\cup_{i=1}^{n}\mathcal{O}_{i}$.
@@
 2), which results in an eroded workspace $\mathcal{W}^{\epsilon}:=\{\mathbf{x}\in\mathbb{R}^{2}\mid\text{d}_{\complement\mathcal{W}^{\ast}}(\mathbf{x})\geq r+\epsilon\}$ and $n$ augmented obstacle regions $\mathcal{O}_{i}^{\epsilon}:=\{\mathbf{x}\in\mathbb{R}^{2}\mid\beta_{i}(\mathbf{x})<\epsilon\}$, $i\in\mathbb{I}$.
-By doing so, the free space of $\bm{P}$ reduces to $\mathcal{X}_{\epsilon}:=\mathcal{W}^{\epsilon}\setminus\mathcal{O}^{\epsilon}$ with $\mathcal{O}^{\epsilon}:=\cup_{i=1}^{n}\mathcal{O}_{i}^{\epsilon}$.
-Figure 2: Schematic diagram of the operating environment.
-The dark gray balls denote the obstacles $\mathcal{O}_{i}^{\ast}$, $i\in\mathbb{I}$, the light gray regions denote the augmented workspace boundary and obstacles, whereas the light blue regions denote the safety margin.
+Then, the free space of $\bm{P}$ reduces to $\mathcal{X}_{\epsilon}:=\mathcal{W}^{\epsilon}\setminus\mathcal{O}^{\epsilon}$ with $\mathcal{O}^{\epsilon}:=\cup_{i=1}^{n}\mathcal{O}_{i}^{\epsilon}$.
+Figure 2: Schematic diagram of the operating environment, where the dark gray balls denote the actual obstacles $\mathcal{O}_{i}^{\ast}$, the light gray regions denote the augmented obstacle regions $\mathcal{O}_{i}$, and the light blue regions denote the safety margin.
+In addition, the dashed circles are the influence regions of obstacles, whereas the pink and yellow lines denote the discontinuous trajectory and its continuous counterpart, respectively.
 II-C Problem formulation
-The collision-free navigation problem for the WMR is now formulated as follows:
+The robot navigation problem is formulated as follows:
 Problem 1.
-[…]
+Consider the WMR kinematics described by (3) under Assumptions 1 and 2.
+The objective is to derive a control law $\mathbf{u}$ that derives the virtual control point $\bm{P}$ from an initial position $\mathbf{x}(0)\in\mathcal{X}_{\epsilon}$ to a designated goal $\mathbf{x}^{\ast}\in\text{int}(\mathcal{X}_{\epsilon})$, without colliding with any obstacles along the route.
 III Main Results
-To solve Problem 1, we propose a closed-form robot navigation scheme that integrates a path planner and a tube-following controller.
-The block diagram of this scheme is depicted in Fig.
-3.
-The path planner explicitly considers the safety margin and generates a collision-free reference trajectory $\mathbf{x}_{d}$ in the free space $\mathcal{X}{\epsilon}$, using tangent cones and the well-known Nagumo’s invariance theorem.
-In the control module, an adaptive tube-following controller is designed to track the planned trajectory in the presence of unknown disturbances, while ensuring that the position $\mathbf{x}$ of $\bm{P}$ stays within a predefined tube around the reference trajectory $\mathbf{x}_{d}$.
-Under this framework, if the radius of the tube is set less than the width $\epsilon$ of the safety margin, then the actual motion trajectory of $\mathbf{x}$ will remain within the set $\mathcal{X}$, thus achieving safe robot motions.
-Figure 3: Block diagram of the collision-free navigation scheme.
+In this section, an approach, called InPTC, is proposed to solve Problem 1.
+The block diagram of this scheme is shown in Fig.
+3, where the planner module generates a collision-free reference trajectory $\mathbf{x}_{d}$ with prescribed-time convergence.
+In the control module, a tube-following controller is derived to achieve prescribed-time trajectory tracking within a safe tube around $\mathbf{x}_{d}$ (with radius less than the safety margin).
+Under this framework, $\mathbf{x}$ will converge to the goal $\mathbf{x}^{\ast}$ within a prescribed time, without colliding with any obstacles along the way, thus achieving prescribed-time safe robot navigation.
+Figure 3: Block diagram of the proposed InPTC scheme.
 III-A Preliminaries
-The definition of tangent cones and Nagumo’s theorem that are necessary for collision-free path planning are given.
 Definition 1 (Bouligand’s tangent cone [28]).
 Given a closed set $\mathcal{F}\in\mathbb{R}^{n}$, the tangent cone to $\mathcal{F}$ at a point $\mathbf{x}\in\mathbb{R}^{n}$ is the subset of $\mathbb{R}^{n}$ defined by
-$\mathbf{T}_{\mathcal{F}}(\mathbf{x}):=\left\{\mathbf{z}\in\mathbb{R}^{n}\mid\mathop{\lim\inf}\limits_{\tau\to 0^{+}}\dfrac{{\rm d}_{\mathcal{F}}(\mathbf{x}+\tau\mathbf{z})}{\tau}=0\right\}.$
+$\mathbf{T}_{\mathcal{F}}(\mathbf{x}):=\left\{\mathbf{z}\in\mathbb{R}^{n}\bigg{|}\mathop{\lim\inf}\limits_{\tau\to 0^{+}}\dfrac{{\rm d}_{\mathcal{F}}(\mathbf{x}+\tau\mathbf{z})}{\tau}=0\right\}.$
 The tangent cone $\mathbf{T}_{\mathcal{F}}(\mathbf{x})$ is a set that contains all the vectors pointing from $\mathbf{x}$ inside or tangent to $\mathcal{F}$, while for $\mathbf{x}\notin\mathcal{F}$, $\mathbf{T}_{\mathcal{F}}(\mathbf{x})=\emptyset$.
@@
 Theorem 1 (Nagumo 1942 [29]).
-Consider the system $\dot{\mathbf{x}}(t)=\mathbf{f}(\mathbf{x}(t))$ and assume that it admits a unique solution in forward time for each initial condition $\mathbf{x}(0)$ in an open set $\mathcal{O}$.
-Then, the closed set $\mathcal{F}\subset\mathcal{O}$ is forward invariant iff
-$\mathbf{f}(\mathbf{x})\in\mathbf{T}_{\mathcal{F}}(\mathbf{x}),~{}\forall\mathbf{x}\in\mathcal{F}.$ (9)
-An intuitive interpretation of Nagumo’s theorem is that if the vector field $\mathbf{f}(\mathbf{x})$ points inside or is tangent to the set $\mathcal{F}$ at each point $\mathbf{x}\in\mathcal{F}$, then $\mathcal{F}$ is forward invariant.
-As stated earlier, $\mathbf{T}_{\mathcal{F}}(\mathbf{x})$ is non-trivial only on $\partial\mathcal{F}$, thus it is only necessary to check (9) for the boundary points.
-Lemma 1.
-Given any constant $b>0$, the following inequalities hold for all $0\leq q<b$:
-$\frac{q}{b}\leq\ln\dfrac{b}{b-q}\leq\dfrac{q}{b-q}.$
-Proof. Define two functions $h_{1}(q):=q/(b-q)-\ln(b/(b-q))$ and $h_{2}(q):=\ln(b/(b-q))-q/b$. After simple calculations, it follows that
-$\dfrac{\partial h_{1}(q)}{\partial q}=\dfrac{q}{(b-q)^{2}}\geq 0,~{}\dfrac{\partial h_{2}(q)}{\partial q}=\dfrac{q}{b(b-q)}\geq 0,$
-[…]
-III-B Path planing based on tangent cones
-Based on Bouligand’s tangent cones and Nagumo’s theorem, a safe path planning approach is proposed, which can guide the virtual control point $\bm{P}$ towards the goal $\mathbf{x}^{\ast}$ (motion-to-goal) from almost [^1] The free space $\mathcal{X}_{\epsilon}$ is a non-contractible space, for which the topological obstruction precludes the possibility of any continuous vector fields achieving global asymptotic stability.
-The basin of attraction of the desired equilibrium must generally exclude at least a set of measure zero.
-all initial conditions $\mathbf{x}(0)\in\mathcal{X}_{\epsilon}$, while ensuring forward invariance of the free space $\mathcal{X}_{\epsilon}$ (collision avoidance).
-We neglect the disturbance $\mathbf{u}_{d}$ in (3) and consider a control law of the following form
-$\mathbf{u}=\mathbf{R}^{-1}(\theta)\bm{\tau},$ (10)
-where $\bm{\tau}\in\mathbb{R}^{2}$ is a virtual control law to be designed. Then, substituting (10) into (3), one gets
+Consider the system $\dot{\mathbf{x}}(t)=\mathbf{f}(\mathbf{x}(t))$, which admits a unique solution in forward time for each initial condition $\mathbf{x}(0)$ in an open set $\mathcal{O}$.
+The closed set $\mathcal{F}\subset\mathcal{O}$ is forward invariant iff $\mathbf{f}(\mathbf{x})\in\mathbf{T}_{\mathcal{F}}(\mathbf{x})$, $\forall\mathbf{x}\in\mathcal{F}$.
+Definition 2 (Time scale transformation function, TSTF [27]).
+A smooth function $\eta:[0,\infty)\to[0,T)$ is called a TSTF if 1) it is strictly increasing; 2) it is s.t.
+$\eta(0)=0$ and $\eta^{\prime}(0)=1$; 3) it is s.t.
+$\lim_{s\to\infty}\eta(s)=T$ and $\lim_{s\to\infty}\eta^{\prime}(s)=0$.
+A TSTF squeezes the infinite-time interval $s\in[0,\infty)$ into a prescribed finite time interval $t:=\eta(s)\in[0,T)$, which plays a key role for achieving prescribed-time planning and control.
+In this work, $\eta(s)$ is of the form
+$\eta(s)=T(1-e^{-\frac{s}{T}}).$ (6)
+III-B Prescribed-time path planning
+We neglect the disturbance $\mathbf{u}_{d}$ in (3) and consider a control law $\mathbf{u}=\mathbf{R}^{-1}(\theta)\bm{\tau}$, where $\bm{\tau}\in\mathbb{R}^{2}$ is a virtual control law to be designed.
+Then, substituting it into (3), one gets
 $\dot{\mathbf{x}}=\bm{\tau}.$ (7)
-According to Nagumo’s theorem, it is sufficient to address the collision-free path planning problem by solving the following constrained optimization problem
-$\begin{split}\min\limits_{\bm{\tau}}&~{}\|\bm{\tau}-\bm{\kappa}_{0}(\mathbf{x})\|,\\ \text{s.t.}&~{}\bm{\tau}\in\mathbf{T}_{\mathcal{{X}}_{\epsilon}}(\mathbf{x}),~{}\forall\mathbf{x}\in\mathcal{X}_{\epsilon},\end{split}$ (12)
-where $\bm{\kappa}_{0}(\mathbf{x})$ is a nominal control law for motion-to-goal and is designed here as the following smooth function:
-$\bm{\kappa}_{0}(\mathbf{x})=-k_{0}(\mathbf{x})(\mathbf{x}-\mathbf{x}^{\ast}),$ (13)
-with $k_{0}(\mathbf{x})>0$ given by
-$k_{0}(\mathbf{x}):=\dfrac{\alpha}{\sqrt{\|\mathbf{x}-\mathbf{x}^{\ast}\|^{2}+\beta^{2}}},$ (14)
-where $\alpha$ and $\beta$ are positive design constants.
-From (13) and (14), it follows that $\|\bm{\kappa}_{0}(\mathbf{x})\|\leq\alpha$, which allows us to generate a saturated nominal vector field.
-The problem (12) is clearly a nearest point problem, which has an explicit solution that will be demonstrated in the following.
-Figure 4: The resulting smooth navigation trajectories maintains a safety margin of size $\epsilon$ (light blue regions) from the workspace boundary and obstacles.
-The dashed circles denote the influence regions from each obstacle.
-Moreover, the orange points are the unstable stationary points associated with each obstacles.
-The robot workspace $\mathcal{W}^{\ast}$ is a convex set, so does $\mathcal{W}^{\epsilon}$, which together with the fact that $\mathbf{x}^{\ast}\in\text{int}(\mathcal{X}_{\epsilon})$ suggests that for all $\mathbf{x}\in\partial\mathcal{W}^{\epsilon}$, $\bm{\kappa}_{0}(\mathbf{x})$ points inside the free space $\mathcal{X}_{\epsilon}$ (i.e., $\bm{\kappa}_{0}(\mathbf{x})\in\mathbf{T}_{\mathcal{{X}}_{\epsilon}}(\mathbf{x})$) and thus is a solution to (12).
-Moreover, for all $\mathbf{x}\in\text{int}(\mathcal{X}_{\epsilon})$, it follows that $\bm{\kappa}_{0}(\mathbf{x})$ is also the solution to (12).
+A tangent cone based control policy (vector field) $\bm{\tau}=\mathbf{h}(\mathbf{x})$ is designed to achieve collision-free path planning. According to Theorem 1, we consider a nearest point problem
+$\begin{split}\min\limits_{\mathbf{h}}&~{}\|\mathbf{h}(\mathbf{x})-\bm{\kappa}_{0}(\mathbf{x})\|\\ \text{s.t.}&~{}\mathbf{h}(\mathbf{x})\in\mathbf{T}_{\mathcal{{X}}_{\epsilon}}(\mathbf{x}),~{}\forall\mathbf{x}\in\mathcal{X}_{\epsilon},\end{split}$ (8)
+where $\bm{\kappa}_{0}(\mathbf{x})$ is a nominal control law for motion-to-goal and is designed here as
+$\bm{\kappa}_{0}(\mathbf{x})=-k_{0}(\mathbf{x}-\mathbf{x}^{\ast}),$ (9)
+with $k_{0}>0$ being a constant.
+The robot workspace $\mathcal{W}^{\ast}$ is a convex set, so does $\mathcal{W}^{\epsilon}$, which suggests that for all $\mathbf{x}\in\partial\mathcal{W}^{\epsilon}$, $\bm{\kappa}_{0}(\mathbf{x})$ points inside the free space $\mathcal{X}_{\epsilon}$ (i.e., $\bm{\kappa}_{0}(\mathbf{x})\in\mathbf{T}_{\mathcal{{X}}_{\epsilon}}(\mathbf{x})$) and thus is a solution to (8).
+Moreover, for all $\mathbf{x}\in\text{int}(\mathcal{X}_{\epsilon})$, $\bm{\kappa}_{0}(\mathbf{x})$ is also the solution to (8).
 Next we check the obstacle boundary points $\mathbf{x}\in\partial\mathcal{O}^{\epsilon}$.
-As $\partial\mathcal{O}^{\epsilon}$ is a smooth hypersurface of $\mathbb{R}^{2}$, which is orientable, there exists a continuously differentiable map (known as the Gauss map [28]) $\mathbf{n}:\partial\mathcal{O}^{\epsilon}\to\mathbb{S}^{1}$ such that for all $\mathbf{x}\in\partial\mathcal{O}^{\epsilon}$, $\mathbf{n}(\mathbf{x})$ is the outward unit normal vector to $\partial\mathcal{O}^{\epsilon}$.
+As $\partial\mathcal{O}^{\epsilon}$ is a smooth hypersurface of $\mathbb{R}^{2}$ and is orientable, there exists a continuously differentiable map (known as Gauss map [30]) $\mathbf{n}:\partial\mathcal{O}^{\epsilon}\to\mathbb{S}^{1}$ such that for all $\mathbf{x}\in\partial\mathcal{O}^{\epsilon}$, $\mathbf{n}(\mathbf{x})$ is the outward unit normal vector to $\partial\mathcal{O}^{\epsilon}$.
 As clearly seen in Fig.
-2, the tangent cone at any $\mathbf{x}\in\partial\mathcal{O}^{\epsilon}$ is a half-space
-$\mathbf{T}_{\mathcal{X}_{\epsilon}}(\mathbf{x}):=\{\mathbf{y}\in\mathbb{R}^{2}\mid(\mathbf{y}-\mathbf{x})^{\top}\mathbf{n}(\mathbf{x})\leq 0\},$ (15)
-which is a convex function.
+2, the tangent cone at any $\mathbf{x}\in\partial\mathcal{O}^{\epsilon}$ is a half-space $\mathbf{T}_{\mathcal{X}_{\epsilon}}(\mathbf{x}):=\{\mathbf{y}\in\mathbb{R}^{2}\mid(\mathbf{y}-\mathbf{x})^{\top}\mathbf{n}(\mathbf{x})\leq 0\}$, which is a convex function.
 Thus, (8) has a unique solution.
-Based on the above argument, it is known that if $\mathbf{x}\in\mathcal{X}_{\epsilon}\setminus\partial\mathcal{O}^{\epsilon}$ or $\mathbf{x}\in\partial\mathcal{O}^{\epsilon}$ and $\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{n}(\mathbf{x})\leq 0$, then $\bm{\kappa}_{0}(\mathbf{x})$ is a solution to (12).
-On the other hand, if $\mathbf{x}\in\mathcal{O}^{\epsilon}$ and $\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{n}(\mathbf{x})>0$, the closest point of (12) is obtained by the orthogonal projection $\mathbf{Q}(\mathbf{n}(\mathbf{x})):=\mathbf{I}_{2}-\mathbf{n}(\mathbf{x})\mathbf{n}^{\top}(\mathbf{x})$ onto the tangent hyperplane of $\partial\mathcal{X}_{\epsilon}$, defined by $\mathbf{H}_{\partial\mathcal{X}_{\epsilon}}(\mathbf{x}):=\{\mathbf{y}\in\mathbb{R}^{2}\mid(\mathbf{y}-\mathbf{x})^{\top}\mathbf{n}(\mathbf{x})=0\}$.
-Then, a general solution to (12) is as follows:
-$\bm{\tau}(\mathbf{x})=\left\{\begin{split}&\bm{\kappa}_{0}(\mathbf{x}),&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})>\epsilon~{}\text{or}\\ &&&\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{n}(\mathbf{x})\leq 0,\\ &\mathbf{Q}(\mathbf{n}(\mathbf{x}))\bm{\kappa}_{0}(\mathbf{x}),&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})=\epsilon~{}\text{and}\\ &&&\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{n}(\mathbf{x})>0.\end{split}\right.$ (16)
-By inspecting (16), we find that the resulting vector field is discontinuous at some boundary points $\mathbf{x}\in\partial\mathcal{O}^{\epsilon}$ like $\bm{P}_{1}-\bm{P}_{4}$ in Fig.
+In summary, if $\mathbf{x}\in\mathcal{X}_{\epsilon}\setminus\partial\mathcal{O}^{\epsilon}$ or $\mathbf{x}\in\partial\mathcal{O}^{\epsilon}\wedge\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{n}(\mathbf{x})\leq 0$, then $\bm{\kappa}_{0}(\mathbf{x})$ is a solution to (8).
+While if $\mathbf{x}\in\mathcal{O}^{\epsilon}\wedge\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{n}(\mathbf{x})>0$, the closest point of (8) is obtained by the orthogonal projection $\mathbf{Q}(\mathbf{n}(\mathbf{x})):=\mathbf{I}_{2}-\mathbf{n}(\mathbf{x})\mathbf{n}^{\top}(\mathbf{x})$ onto the tangent hyperplane of $\partial\mathcal{X}_{\epsilon}$, defined by $\mathbf{H}_{\partial\mathcal{X}_{\epsilon}}(\mathbf{x}):=\{\mathbf{y}\in\mathbb{R}^{2}\mid(\mathbf{y}-\mathbf{x})^{\top}\mathbf{n}(\mathbf{x})=0\}$.
+Then, a general solution to (8) is as follows: $\mathbf{h}(\mathbf{x})=\bm{\kappa}_{0}(\mathbf{x})$ if $\text{d}_{\mathcal{O}}(\mathbf{x})>\epsilon$ or $\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{n}(\mathbf{x})\leq 0$, and $\mathbf{h}(\mathbf{x})=\mathbf{Q}(\mathbf{n}(\mathbf{x}))\bm{\kappa}_{0}(\mathbf{x})$ if $\text{d}_{\mathcal{O}}(\mathbf{x})=\epsilon\wedge\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{n}(\mathbf{x})>0$.
+The resulting vector field is discontinuous at some boundary points $\mathbf{x}\in\partial\mathcal{O}^{\epsilon}$ like the green points in Fig.
 2.
-To address this issue, a continuous version of (16) is proposed in the sequel.
+To address this issue, a continuous control law is proposed in the sequel.
 Following the line of [4], we specify an influence region for each obstacle (marked by the dashed line in Fig.
-4), which is defined as $\{\mathbf{x}\in\mathbb{R}^{2}\mid\text{d}_{\mathcal{O}_{i}}(\mathbf{x})\leq\epsilon^{\ast}\}$, where $\epsilon^{\ast}\in(\epsilon,h]$ and $i\in\mathbb{I}$.
+2), which is defined as $\{\mathbf{x}\in\mathbb{R}^{2}\mid\text{d}_{\mathcal{O}_{i}}(\mathbf{x})\leq\epsilon^{\ast}\}$, where $\epsilon^{\ast}\in(\epsilon,h]$ and $i\in\mathbb{I}$.
 Obstacle avoidance is activated only when the position $\mathbf{x}$ of $\bm{P}$ enters the influence regions of obstacles.
 To proceed, a bearing vector is defined as
-$\mathbf{b}(\mathbf{x}):=\dfrac{\mathbf{P}_{\partial\mathcal{O}}(\mathbf{x})-\mathbf{x}}{\|\mathbf{P}_{\partial\mathcal{O}}(\mathbf{x})-\mathbf{x}\|}=\dfrac{\mathbf{P}_{\partial\mathcal{O}}(\mathbf{x})-\mathbf{x}}{\text{d}_{\mathcal{O}}(\mathbf{x})}.$ (17)
-for all $\mathbf{x}$ with $\epsilon\leq\text{d}_{\mathcal{O}}(\mathbf{x})\leq\epsilon^{\ast}$.
-Since $\text{d}(\mathcal{O}_{i}^{\ast},\mathcal{O}_{j}^{\ast})>2(r+h)$, $\forall i,j\in\mathbb{I}$, $i\neq j$ (referring to (5) in Assumption 2) and $\epsilon^{\ast}\leq h$, there can be only one obstacle $\mathcal{O}_{i}$, $i\in\mathbb{I}$ such that $\text{d}_{\mathcal{O}}(\mathbf{x})=\text{d}_{\mathcal{O}_{i}}(\mathbf{x})\leq\epsilon^{\ast}$.
+$\mathbf{b}(\mathbf{x}):=\dfrac{\mathbf{P}_{\partial\mathcal{O}}(\mathbf{x})-\mathbf{x}}{\|\mathbf{P}_{\partial\mathcal{O}}(\mathbf{x})-\mathbf{x}\|}=\dfrac{\mathbf{P}_{\partial\mathcal{O}}(\mathbf{x})-\mathbf{x}}{\text{d}_{\mathcal{O}}(\mathbf{x})},$ (10)
+where $\mathbf{P}_{\partial\mathcal{O}}(\mathbf{x}):=\{\mathbf{y}\in\partial\mathcal{O}\mid\|\mathbf{y}-\mathbf{x}\|=\text{d}_{\mathcal{O}}(\mathbf{x})\}$ is a set-valued map.
+Since $\text{d}(\mathcal{O}_{i}^{\ast},\mathcal{O}_{j}^{\ast})>2(r+h)$, $\forall i,j\in\mathbb{I}$, $i\neq j$ (referring to Assumption 2) and $\epsilon^{\ast}\leq h$, there can be only one obstacle $\mathcal{O}_{i}$ such that $\text{d}_{\mathcal{O}}(\mathbf{x})=\text{d}_{\mathcal{O}_{i}}(\mathbf{x})\leq\epsilon^{\ast}$.
 With this in mind, $\mathbf{b}(\mathbf{x})$ in (10) can be computed by $\mathbf{b}(\mathbf{x})=(\mathbf{c}_{i}-\mathbf{x})/\|\mathbf{c}_{i}-\mathbf{x}\|$.
 Note that when $\mathbf{x}\in\partial\mathcal{O}_{\epsilon}$, $\mathbf{b}(\mathbf{x})$ is equivalent to the Gauss map $\mathbf{n}(\mathbf{x})$.
-Now the control law (16) is modified as
-$\bm{\tau}(\mathbf{x})=\left\{\begin{split}&\bm{\kappa}_{0}(\mathbf{x}),&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})>\epsilon^{\ast}~{}\text{or}\\ &&&\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})\leq 0,\\ &\mathbf{\Pi}(\mathbf{x})\bm{\kappa}_{0}(\mathbf{x}),&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})\leq\epsilon^{\ast}~{}\text{and}\\ &&&\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})>0,\end{split}\right.$ (18)
+Now the control law is modified as
+$\mathbf{h}(\mathbf{x})=\left\{\begin{split}&\bm{\kappa}_{0}(\mathbf{x}),&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})>\epsilon^{\ast}~{}\text{or}\\ &&&\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})\leq 0,\\ &\mathbf{\Pi}(\mathbf{x})\bm{\kappa}_{0}(\mathbf{x}),&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})\leq\epsilon^{\ast}~{}\text{and}\\ &&&\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})>0,\end{split}\right.$ (11)
 with
 $\mathbf{\Pi}(\mathbf{x}):=\mathbf{I}_{2}-\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))\mathbf{b}(\mathbf{x})\mathbf{b}^{\top}(\mathbf{x}),$ (12)
-where $\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))\in\mathbb{R}$ is a $C^{1}$ bump function that smoothly transitions from $1$ to 00 on the interval $\text{d}_{\mathcal{O}}(\mathbf{x})\in[\epsilon,\epsilon^{\ast}]$, that is to say, its first derivatives equal to zero at the endpoints of the interval $[\epsilon,\epsilon^{\ast}]$.
+where $\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))\in\mathbb{R}$ is a $C^{1}$ bump function that smoothly transitions from $1$ to 00 on the interval $\text{d}_{\mathcal{O}}(\mathbf{x})\in[\epsilon,\epsilon^{\ast}]$.
 A simple choice of $\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))$ is
-$\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))=\left\{\begin{split}&1,&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})\leq\epsilon,\\ &\lambda(\text{d}_{\mathcal{O}}(\mathbf{x})),&\text{if}~{}&\epsilon<\text{d}_{\mathcal{O}}(\mathbf{x})<\epsilon^{\ast},\\ &0,&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})\geq\epsilon^{\ast},\end{split}\right.$ (20)
-with
-$\lambda(\text{d}_{\mathcal{O}}(\mathbf{x}))=\dfrac{1}{2}\left[1-\cos\left(\pi\dfrac{\epsilon^{\ast}-\text{d}_{\mathcal{O}}(\mathbf{x})}{\epsilon^{\ast}-\epsilon}\right)\right].$ (21)
-One can verify that $\phi(\epsilon)=1$, $\phi(\epsilon^{\ast})=0$, and
-$\left.\dfrac{d\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))}{d\text{d}_{\mathcal{O}}(\mathbf{x})}\right|_{\text{d}_{\mathcal{O}}(\mathbf{x})=0~{}\text{or}~{}\epsilon^{\ast}}=0.$
-Since the nominal control law, $\bm{\kappa}_{0}(\mathbf{x})$, the bearing vector, $\mathbf{b}(\mathbf{x})$, and the bump function, $\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))$, are continuously differentiable, the resulting control law $\bm{\tau}(\mathbf{x})$ in (18) is continuous and piecewise continuously differentiable on the domain $\mathcal{X}$.
-It is not differentiable only at those points satisfying $\text{d}_{\mathcal{O}}(\mathbf{x})\leq\epsilon^{\ast}$ and $\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})=0$.
-Now the main result of the path planning problem is summarized in the following theorem.
+$\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))=\left\{\begin{aligned} &1,&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})\leq\epsilon,\\ &\lambda(\text{d}_{\mathcal{O}}(\mathbf{x})),&\text{if}~{}&\epsilon<\text{d}_{\mathcal{O}}(\mathbf{x})<\epsilon^{\ast},\\ &0,&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})\geq\epsilon^{\ast},\end{aligned}\right.$ (13)
+where $\lambda(\text{d}_{\mathcal{O}}(\mathbf{x}))=0.5[1-\cos(\pi(\epsilon^{\ast}-\text{d}_{\mathcal{O}}(\mathbf{x}))/(\epsilon^{\ast}-\epsilon))]$.
+It is clear that the modified control law (11) is continuous and piecewise continuously differentiable on the domain $\mathcal{X}$.
 Theorem 2.
-[…]
+Consider the kinematics defined by (7).
+If $\mathbf{x}^{\ast}\in\text{int}(\mathcal{X}_{\epsilon})$ and the obstacles $\mathcal{O}_{i}^{\ast}$, $i\in\mathbb{I}$ satisfying Assumption 2, then the continuous control law (11) with $\bm{\kappa}_{0}(\mathbf{x})$ given by (9) can guarantee that:
   1. 1.
-    The free space $\mathcal{X}_{\epsilon}$ in (7) is forward invariant.
+    The free space $\mathcal{X}_{\epsilon}$ is forward invariant.
   2. 2.
-    For any $\mathbf{x}(0)\in\mathcal{X}_{\epsilon}$, the kinematics (11) admits a unique solution in forward time, which asymptotically converges to the set $\{\mathbf{x}^{\ast}\}\cup_{i=1}^{n}\mathcal{S}_{i}$ with $\mathcal{S}_{i}$ given by
-    $\mathcal{S}_{i}:=\left\{\mathbf{x}\in\mathcal{X}_{\epsilon}\bigg{|}{\rm d}_{\mathcal{O}_{i}}(\mathbf{x})=\epsilon,~{}\dfrac{\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})}{\|\bm{\kappa}_{0}(\mathbf{x})\|}=1\right\}.$ (22)
+    For any $\mathbf{x}(0)\in\mathcal{X}_{\epsilon}$, the solution of (7) admits a unique solution in forward time, which asymptotically converges to the set $\mathcal{C}:=\{\mathbf{x}^{\ast}\}\cup_{i=1}^{n}\{\mathbf{s}_{i}\}$, where the stationary point $\mathbf{s}_{i}$ that satisfies ${\rm d}_{\mathcal{O}_{i}}(\mathbf{s}_{i})=\epsilon\wedge\bm{\kappa}_{0}^{\top}(\mathbf{s}_{i})\mathbf{b}(\mathbf{s}_{i})/\|\bm{\kappa}_{0}(\mathbf{s}_{i})\|=1$ is locally unstable.
   3. 3.
-    The undesired stationary points $\mathbf{s}_{i}\in\mathcal{S}_{i}$, associated with obstacles $\mathcal{O}_{i}$, are locally unstable and only have a stable manifold of measure zero.
-  4. 4.
     The equilibrium point $\mathbf{x}=\mathbf{x}^{\ast}$ is almost globally asymptotically stable and locally exponentially stable.
 […]
+To further achieve prescribed-time path planning, the TSTF $t:=\eta(s)$ defined in (6) is introduced to squeeze and map the collision-free trajectory of (7) from an infinite time interval $s\in[0,\infty)$ to a prescribed finite time interval $t\in[0,T)$.
+Let $\eta^{\prime}(s):=d\eta(s)/ds$ and define a prescribed-time gain function (PTGF) $\alpha:\mathbb{R}_{\geq 0}\to\mathbb{R}_{>0}$ as
+$\alpha(t)=\alpha(\eta(s)):=\dfrac{1}{\eta^{\prime}(s)}=\dfrac{T}{T-t},~{}t\in[0,T),$ (14)
+showing that $\alpha(t)>0$ is continuously differentiable on $t\in[0,T)$ and satisfies $\alpha(0)=1$ and $\lim_{t\to T}\alpha(t)=\infty$.
+Here the prescribed-time control law is designed as $\bm{\tau}=\mathbf{h}^{\ast}(\mathbf{x},t)$, where $\mathbf{h}^{\ast}$ is given by
+$\mathbf{h}^{\ast}(\mathbf{x},t)=\left\{\begin{split}&\alpha(t)\mathbf{h}(\mathbf{x}),&t&\in[0,T),\\ &\mathbf{h}(\mathbf{x}),&t&\in[T,\infty),\end{split}\right.$ (15)
+with $\mathbf{h}(\mathbf{x})$ and $\alpha(t)$ defined in (11) and (14), respectively.
+Theorem 3.
+Consider the kinematics given by (7) with initial conditions satisfying $\mathbf{x}(0)\in\mathcal{X}_{\epsilon}$.
+Then, the closed-loop system $\dot{\mathbf{x}}=\mathbf{h}^{\ast}(\mathbf{x},t)$ converge to the set $\mathcal{C}$ at $t=T$, while avoiding obstacle regions $\mathcal{O}_{i}^{\epsilon}$, $i\in\mathbb{I}$.
+The system under $\bm{\tau}=\mathbf{h}^{\ast}(\mathbf{x},t)$ follows the same path as the closed-loop system $\dot{\mathbf{x}}=\mathbf{h}(\mathbf{x})$, but goes faster within the time window $[0,T)$.
+Proof: The proof is conducted over two time intervals.
+1) Consider the interval $t\in[0,T)$. The closed-loop system is written as
+$\dot{\mathbf{x}}(t)=\alpha(t)\mathbf{h}(\mathbf{x}(t)).$ (16)
+Let $\bar{\mathbf{x}}(s):=\mathbf{x}(t)=\mathbf{x}(\eta(s))$. Then, the system (16) is rewritten in the stretched infinite-time interval $s\in[0,\infty)$ as
+$\bar{\mathbf{x}}^{\prime}(s):=\dfrac{d\bar{\mathbf{x}}(s)}{ds}=\dfrac{d\mathbf{x}(\eta(s))}{d\eta(s)}\cdot\dfrac{d\eta(s)}{ds}.$ (17)
+In view of 14 and (16), 17 becomes
+$\bar{\mathbf{x}}^{\prime}(s)=\mathbf{h}(\bar{\mathbf{x}}(s)).$ (18)
+The initial conditions satisfy $\bar{\mathbf{x}}^{\prime}(0)=\dot{\mathbf{x}}(0)$ and $\bar{\mathbf{x}}(0)=\mathbf{x}(0)$, where the facts that $\eta(0)=0$ and $\alpha(0)=1$ have been used.
+As (18) has the same solution as the closed-loop system $\dot{\mathbf{x}}(t)=\mathbf{h}(\mathbf{x}(t))$, it follows from Theorem 2 that $\bar{\mathbf{x}}(s)$ asymptotically converges to the set $\mathcal{C}$.
+This implies that $\mathbf{x}(t)$ converges to the set $\mathcal{C}$ at the prescribed time $T$, due to $\mathbf{x}(t)=\bar{\mathbf{x}}(s)$ and $t\to T$ as $s\to\infty$.
+Additionally, Theorem 2 ensures that $\bar{\mathbf{x}}(s)$ is collision-free on $t\in[0,\infty)$.
+As a result, $\mathbf{x}(t)$ is also collision-free on $t\in[0,T)$.
+[…]
 Remark 2.
-The proposed safe path planning scheme offers an analytical solution and thus is computationally much more efficient than the optimization-based methods.
-Moreover, unlike the navigation functions, the scheme can achieve almost global asymptotic convergence without requiring unknown parameter tuning, making it easy to implement in practice.
+The proposed prescribed-time control policy (15) is discontinuous at $t=T$ and cannot be practically implemented due to the unboundedness of $\alpha(t)$ as $t\to T$.
+To remedy this, a continuous version is given as
+$\tau=\mathbf{h}^{\ast}(\mathbf{x},t)=\alpha^{\ast}(t)\mathbf{h}(\mathbf{x}),$ (19)
+with the PTGF given by
+$\alpha^{\ast}(t):=\left\{\begin{aligned} &\alpha(t),&~{}&t\in[0,T^{\ast}),\\ &\alpha(T^{\ast}),&~{}&t\in[T^{\ast},\infty),\end{aligned}\right.$ (20)
+where $T^{\ast}:=T-\varsigma>0$, and $0<\varsigma\ll T$ is a sufficiently small constant.
+The resulting continuous vector field can guide $\mathbf{x}$ to a small neighborhood of the set $\mathcal{C}$ at $t=T^{\ast}$.
+Since $\alpha^{\ast}(t)\equiv T/\varsigma>0$ on $[T^{\ast},\infty)$, $\mathbf{x}(t)$ will asymptotically converge to the set $\mathcal{C}$ after $t=T^{\ast}$.
+It is clear from (9) that choosing a larger $k_{0}$ will increase the convergence rate of the planner during the transient-state phase.
+However, this comes at the expense of increased reference velocity.
 Remark 3.
-A smooth and saturated nominal control law (18) is proposed, instead of the traditional proportional control law $k_{0}(\mathbf{x})=-k(\mathbf{x}-\mathbf{x}^{\ast})$ [4, 21], with the aim of generating a saturated vector field (18).
-This is justified by the inequality $\|\bm{\tau}(\mathbf{x})\|\leq\|\bm{\kappa}_{0}(\mathbf{x})\|\leq\alpha$, where the equality $\|\bm{\Pi}(\mathbf{x})\|=1$ has been used.
-In fact, different types of saturation functions can be used to design $\kappa_{0}(\mathbf{x})$ as an alternative, such as the standard $\text{sat}(\cdot)$, hyperbolic tangent function $\tanh(\cdot)$, $\text{arctan}(\cdot)$, etc.
-However, it is worth noting that these functions may result in the nominal control direction not aligning directly with the desired position, which in turn leads to an unnecessarily long path.
-III-C Adaptive tube-following control
-The planning algorithm introduced in Sec.
-III-B can generate a collision-free reference trajectory for the WMR.
-To avoid conflicts of notations, the reference trajectory is denoted as $\mathbf{x}_{d}$, which is governed by
-$\dot{\mathbf{x}}_{d}=\bm{\tau}_{d},$ (23)
-where $\bm{\tau}_{d}$ is given by (18) but with $\mathbf{x}$ replaced by $\mathbf{x}_{d}$.
-Let us define the position tracking error as $\mathbf{x}_{e}:=\mathbf{x}-\mathbf{x}_{d}$ and impose an inequality constraint on it as follows:
+Although the prescribed-time path planner (19) is partially inspired by [4], there remain substantial differences in the following two aspects:
+  1. 1.
+    A $C^{1}$ bump function $\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))$ defined in (13) is introduced into the projection function $\mathbf{\Pi}(\mathbf{x})$, which makes the resulting vector filed more smoother;
+  2. 2.
+    By incorporating a PTGF $\alpha(t)$ (generated from the TSTF defined in (6)) into the planning algorithm developed in [4], the proposed planner (19) achieves prescribed-time collision-free path planning.
+Moreover, the planning algorithm offers an analytical solution without involving any numerical optimization solving, rendering it computationally efficient.
+Remark 4.
+Although the path planner in (19) is proposed for environments with circular obstacles, it can be easily extended to deal with convex polygonal obstacles, such as rectangle, trapezoid, and triangle.
+For such obstacles, we can define the augmented obstacle regions $\mathcal{O}_{i}:=\mathcal{O}_{i}^{\ast}\cup\{\mathbf{x}\in\mathbb{R}^{2}\mid{\rm d}_{\mathcal{O}_{i}^{\ast}}(\mathbf{x})<r\}$ and $\mathcal{O}_{i}^{\epsilon}:=\mathcal{O}_{i}^{\ast}\cup\{\mathbf{x}\in\mathbb{R}^{2}\mid{\rm d}_{\mathcal{O}_{i}^{\ast}}(\mathbf{x})<r+\epsilon\}$.
+The distance function ensures that the boundaries of $\mathcal{O}_{i}$ and $\mathcal{O}_{i}^{\epsilon}$ are smooth, as shown in Fig.
+4.
+In practical implementation, the distance function ${\rm d}_{\mathcal{O}}(\mathbf{x})$ and the bearing vector $\mathbf{b}(\mathbf{x})$ in $\mathbf{h}(\mathbf{x})$ can be obtained locally using on-board LiDARs.
+Figure 4: Augmented regions of convex polygonal obstacles.
+III-C Prescribed-time tube-following control
+To avoid conflicts of notations, the reference trajectory is denoted as $\mathbf{x}_{d}$, which is governed by $\dot{\mathbf{x}}_{d}=\bm{\tau}_{d}$, where $\bm{\tau}_{d}$ is given by (19) but with $\mathbf{x}$ replaced by $\mathbf{x}_{d}$.
+Define the position tracking error as $\mathbf{x}_{e}:=\mathbf{x}-\mathbf{x}_{d}$.
+It follows from (3) that
+$\dot{\mathbf{x}}_{e}=\mathbf{R}(\theta)(\mathbf{u}+\mathbf{u}_{d})-\bm{\tau}_{d}.$ (21)
+To ensure the tracking safety, we impose the following “tube” constraint on $\mathbf{x}_{e}$
 $\|\mathbf{x}_{e}(t)\|^{2}<\rho^{2},$ (22)
@@
 5.
-Since a safety margin is considered in the planning module, if the tube radius is taken no larger than the size $\epsilon$ of the safety margin and (24) holds for all $t\geq 0$, then the actual trajectory $\mathbf{x}(t)$ will remain within the predefined tube without colliding with any actual obstacles, that is, $\mathbf{x}\in\mathcal{X}$.
-Figure 5: Illustration of the predefined tube.
-In the following, an adaptive tube-following controller is developed to achieve high-precision trajectory tracking, while ensuring that $\mathbf{x}(t)$ evolves within the tube defined by (24).
+Since a safety margin is considered in the planning module, if the tube radius is taken no larger than the size $\epsilon$ of the safety margin and (22) holds for all $t\geq 0$, then the actual trajectory $\mathbf{x}(t)$ remains within the safe tube without colliding with any obstacles, that is, $\mathbf{x}(t)\in\mathcal{X}$, $\forall t\geq 0$.
+Figure 5: Illustration of the predefined safe tube.
+In the following, a tube-following controller is developed to achieve prescribed-time trajectory tracking, while ensuring that $\mathbf{x}(t)$ evolves within the tube defined by (22).
 To this end, we define a transformed error
@@
 Inspecting (23) reveals that $10\leq\xi(t)<$ 1 is equivalent to (22), and $\xi(t)=0$ only when $\mathbf{x}_{e}(t)=\mathbf{0}$.
-Therefore, the tube-following control problem boils down to achieving $\lim_{t\to\infty}\xi(t)=0$, while ensuring $10\leq\xi(t)<$ 1, $\forall t\geq 0$, via a properly-designed controller.
-To achieve this, the following logarithmic barrier function is considered
-$L=\dfrac{1}{2}\ln\dfrac{1}{1-\xi}.$ (27)
-Let $\mathbf{z}:=\mathbf{x}_{e}/(\rho^{2}(1-\xi))$. Then, evaluating the time derivative of $L$ along (26) yields
-$\dot{L}=$ $~{}\mathbf{z}^{\top}[\mathbf{R}(\theta)(\mathbf{u}+\mathbf{u}_{d})-\bm{\tau}_{d}]$
-$\leq$ $~{}\mathbf{z}^{\top}(\mathbf{R}(\theta)\mathbf{u}-\bm{\tau}_{d})+d\|\mathbf{z}\|,$ (28)
-wherein we have used the fact that $\|\mathbf{R}(\theta)\|=1$. The adaptive control law is designed as
-$\mathbf{u}=\mathbf{R}^{-1}(\theta)(-k\mathbf{x}_{e}+\bm{\tau}_{d}-\bm{\varpi}),$ (29)
+Therefore, the prescribed-time tube-following control problem boils down to achieving $\lim_{t\to T_{f}}\xi(t)=0$, while ensuring $10\leq\xi(t)<$ 1, $\forall t\geq 0$, via a properly-designed controller.
+To achieve trajectory tracking within a prescribed finite time $T_{f}>0$, as per Definition 2, we introduce a TSTF $t:=\eta_{f}(s)=T_{f}(1-e^{-s/T_{f}})$, which generates the following PTGF:
+$\alpha_{f}(t)=\alpha_{f}(\eta_{f}(s)):=\dfrac{1}{\eta_{f}^{\prime}(s)}=\dfrac{T_{f}}{T_{f}-t},~{}t\in[0,T_{f}).$ (25)
+Let us define $\mathbf{z}:=\mathbf{x}_{e}/(\rho^{2}(1-\xi))$ and design the prescribed-time control law as
+$\mathbf{u}=\alpha_{f}^{\ast}(t)\mathbf{u}_{n}+\mathbf{R}^{-1}(\theta)(1-\alpha_{f}^{\ast}(t))\bm{\tau}_{d},$ (26)
 with
-$\bm{\varpi}:=\dfrac{\hat{d}^{2}\mathbf{z}}{\sqrt{\hat{d}^{2}\|\mathbf{z}\|^{2}+\varphi^{2}}},$ (30)
-where $\hat{d}$ denotes the estimate of $d$ (see Assumption 1 for its definition) and is updated by a projection-based adaptive law.
-Suppose $d\in\Omega:=\{\nu\in\mathbb{R}\mid 0\leq\nu\leq d_{m}\}$ and let $\Omega_{\delta}:=\{\nu\in\mathbb{R}\mid 0\leq\nu\leq d_{m}+\delta\}$, where $d_{m}>0$ and $\delta>0$ are known constants.
-Inspired by [29], the adaptive law is derived as
-$\dot{\hat{d}}=\text{Proj}(\hat{d},\Phi),~{}\Phi:=\|\mathbf{z}\|-\gamma\hat{d},$ (31)
-with
-$\text{Proj}(\hat{d},\Phi)=\left\{\begin{split}&\eta\Phi,&{\rm if}~{}&1)~{}0\leq\hat{d}<d_{m}~{}\text{or}\\ &&~{}&2)~{}\hat{d}\geq d_{m}~{}\text{and}~{}\Phi\leq 0,\\ &\eta\breve{\Phi},&{\rm if}~{}&3)~{}\hat{d}\geq d_{m}~{}\text{and}~{}\Phi>0,\end{split}\right.$ (32)
-where $\eta,\,\gamma>0$ are design constants and
-$\breve{\Phi}=\left(1-\dfrac{\hat{d}-d_{m}}{\delta}\right)\Phi.$ (33)
-Note that the projection operator $\text{Proj}(\hat{d},\Phi)$ is locally Lipschitz in $(\hat{d},\Phi)$ [29].
-Choosing the initial condition of (31) as $\hat{d}(0)\in\Omega_{\delta}$, we will now demonstrate that
-$\hat{d}(0)\in\Omega_{\delta}\Rightarrow\hat{d}(t)\in\Omega_{\delta},~{}\forall t\geq 0.$ (34)
-In Cases 1 and 2, the adaptive law (31) reduces to
-$\dot{\hat{d}}=\eta(\|\mathbf{z}\|-\gamma\hat{d}).$ (35)
-Solving (35) yields
-[…]
-where $t_{0}\geq 0$ marks the beginning of each time interval during which Cases 1 and 2 hold.
-Moreover, in these two cases, (34) holds as a result of the given conditions.
-In Case 3, however, (31) becomes
-$\dot{\hat{d}}=\dfrac{\delta+d_{m}-\hat{d}}{\delta}\Phi,$ (37)
-which implies that $\hat{d}$ decreases when $\hat{d}>\delta+d_{m}$ and increases if $\hat{d}<\delta+d_{m}$ until it reaches $\delta+d_{m}$ such that $\dot{\hat{d}}=0$.
-Therefore, the adaptive law (31) ensures that (34) holds.
-Let us define the parameter estimation error as $\tilde{d}:=\hat{d}-d$.
+$\mathbf{u}_{n}=\mathbf{R}^{-1}(\theta)\left(-k_{1}\mathbf{x}_{e}-\dfrac{k_{2}}{\alpha_{f}^{\ast}(t)}\mathbf{z}+\bm{\tau}_{d}\right),$ (27)
+where $k_{1},k_{2}>0$ are constant gains, and $\alpha_{f}^{\ast}(t)$ is a continuous PTGF akin to $\alpha^{\ast}(t)$ in (20), given by
+$\alpha_{f}^{\ast}(t):=\left\{\begin{aligned} &\alpha_{f}(t),&~{}&t\in[0,T_{f}^{\ast}),\\ &\alpha_{f}(T_{f}^{\ast}),&~{}&t\in[T_{f}^{\ast},\infty),\end{aligned}\right.$ (28)
+with $T_{f}^{\ast}:=T_{f}-\varsigma_{f}>0$ and $0<\varsigma_{f}\ll T_{f}$ being constants.
 Theorem 4.
-Consider the WMR kinematics (3) under Assumption 1 and suppose that $d\in\Omega$.
-Given the initial position $\mathbf{x}(0)$ satisfying (25), then the adaptive controller (29), together with the projection-based update law (31), can steer $\mathbf{x}$ to track the reference trajectory $\mathbf{x}_{d}$ given by (23) along with the predefined tube (25), while ensuring uniform ultimate boundedness of the position error $\mathbf{x}_{e}$ and the estimation error $\tilde{d}$.
-Proof. Consider the Lyapunov function candidate (LFC):
-$V=L+\dfrac{1}{2\eta}\tilde{d}^{2}.$ (38)
-Taking the derivative of $V$ and using (III-C) and (29) lead to
-$\dot{V}\leq-k\mathbf{z}^{\top}\mathbf{x}_{e}-\mathbf{z}^{\top}\bm{\varpi}+\hat{d}\|\mathbf{z}\|+\dfrac{1}{\eta}\tilde{d}(\dot{\hat{d}}-\eta\|\mathbf{z}\|).$ (39)
-Invoking the inequality $|\nu|\leq\nu^{2}/\sqrt{\nu^{2}+\varphi^{2}}+\varphi$ for any $\nu\in\mathbb{R}$ and $\varphi>0$, one gets $-\mathbf{z}^{\top}\bm{\varpi}+\hat{d}\|\mathbf{z}\|\leq\varphi$.
-In addition, noting (26), (III-C), and Lemma 1, it follows that $-k\mathbf{z}^{\top}\mathbf{x}_{e}=-k\xi/(1-\xi)\leq-2kL$ for all $10\leq\xi<$ 1.
+Consider the WMR kinematics (3) under Assumption 1.
+If the initial position $\mathbf{x}(0)$ satisfies (22) and the gain $k_{1}$ is chosen such that $k_{1}T_{f}\gg 1$ holds, then the control law (26) achieves practical prescribed-time convergence of the position error $\mathbf{x}_{e}$, while ensuring it always remains within the safe tube defined by (22).
+Proof. We analyze the closed-loop behavior over two time intervals $[0,T_{f}^{\ast})$ and $[T_{f}^{\ast},\infty)$, separately.
+1) We first consider the time interval $t\in[0,T_{f}^{\ast})$. Substituting (26) into (21) yields
+$\dot{\mathbf{x}}_{e}=-k_{1}\alpha_{f}(t)\mathbf{x}_{e}-k_{2}\mathbf{z}+\mathbf{R}(\theta)\mathbf{u}_{d}.$ (29)
+To proceed, let us define
+$\bar{\mathbf{x}}_{e}(s)$ $:=\mathbf{x}_{e}(\eta_{f}(s)),$ (30a)
+$\bar{\xi}(s)$ $:=\xi(\eta_{f}(s))=\|\bar{\mathbf{x}}_{e}(s)\|^{2}/\rho^{2},$ (30b)
+$\bar{\mathbf{z}}(s)$ $:=\mathbf{z}(\eta_{f}(s))=\bar{\mathbf{x}}_{e}(s)/(\rho^{2}(1-\bar{\xi}(s))).$ (30c)
+Further define $\bar{\mathbf{d}}(s):=\mathbf{R}(\theta(\eta_{f}(s)))\mathbf{u}_{d}(\eta_{f}(s))$ for notational concision.
+Then, the closed-loop system (29) can be rewritten in the stretched infinite time interval $s\in[0,\infty)$ as
+$\bar{\mathbf{x}}_{e}^{\prime}(s)$ $=\dfrac{d\bar{\mathbf{x}}_{e}(s)}{ds}=\dfrac{d\mathbf{x}_{e}(\eta_{f}(s))}{d\eta_{f}(s)}\cdot\dfrac{d\eta_{f}(s)}{ds}$
+$=-k_{1}\bar{\mathbf{x}}_{e}(s)-\dfrac{1}{\alpha_{f}(\eta_{f}(s))}(k_{2}\bar{\mathbf{z}}(s)-\bar{\mathbf{d}}(s)),$ (31)
+with $\bar{\mathbf{x}}_{e}(0)=\mathbf{x}_{e}(0)$ and $\bar{\mathbf{x}}_{e}^{\prime}(0)=\dot{\mathbf{x}}_{e}(0)$.
+Consider the following logarithmic barrier function
+$L(\bar{\xi})=\dfrac{1}{2}\ln\dfrac{1}{1-\bar{\xi}}.$ (32)
+Taking the derivative of $L$ w.r.t. $s$ along (III-C) yields
+$L^{\prime}(\bar{\xi}(s))=\bar{\mathbf{z}}^{\top}(s)\left(-k_{1}\bar{\mathbf{x}}_{e}(s)-\dfrac{k_{2}\bar{\mathbf{z}}(s)-\bar{\mathbf{d}}(s)}{\alpha_{f}(\eta_{f}(s))}\right).$ (33)
+From [31, Lemma 3], it follows that $-k_{1}\bar{\mathbf{z}}^{\top}\bar{\mathbf{x}}_{e}=-k_{1}\bar{\xi}/(1-\bar{\xi})\leq-k_{1}L$ holds for all $10\leq\bar{\xi}<$ 1.
+In addition, recalling Assumption 1 and the fact that $\|\mathbf{R}(\theta)\|=1$, we have $\bar{\mathbf{z}}^{\top}\bar{\mathbf{d}}\leq k_{2}\|\bar{\mathbf{z}}\|^{2}+d/(4k_{2})$.
 As a result, (33) reduces to
-$\dot{V}\leq-2kL+\varphi+\dfrac{1}{\eta}\tilde{d}(\dot{\hat{d}}-\eta\|\mathbf{z}\|),$ (40)
-on the set $10\leq\xi<$ 1. To proceed, we consider three cases in (32). For Cases 1 and 2, it follows from (35) that
-$\tilde{d}(\dot{\hat{d}}-\eta\|\mathbf{z}\|)=-\eta\gamma\tilde{d}\hat{d}.$ (41)
-While for Case 3, we have
-$\tilde{d}(\dot{\hat{d}}-\eta\|\mathbf{z}\|)=-\eta\gamma\tilde{d}\hat{d}-\dfrac{\eta\tilde{d}(\hat{d}-d_{m})\Phi}{\delta}\leq-\eta\gamma\tilde{d}\hat{d},$ (42)
-which is true because $\tilde{d}\geq 0$, $\hat{d}-d_{m}\geq 0$, and $\Phi>0$ according to the conditions specified in Case 3. In view of (41) and (42), one can easily get
-$\dot{V}$ $\leq-2kL-\gamma\tilde{d}\hat{d}+\varphi$
-$\leq-2kL-\dfrac{\gamma}{2}\tilde{d}^{2}+c$
-$\leq-aV+c,$ (43)
-where $a:=\min\{2k,\eta\gamma\}>0$ and $c:=\gamma d^{2}/2+\varphi>0$. Integrating both sides of (III-C) over $[0,\infty)$ yields
-$V(t)\leq\left(V(0)-\dfrac{c}{a}\right)e^{-at}+\dfrac{c}{a},~{}\forall t\geq 0,$ (44)
-showing that $V\in\mathcal{L}_{\infty}$ and hence $L\in\mathcal{L}_{\infty}$.
-The later implies that $10\leq\xi(t)<$ 1 for all $t\geq 0$.
-As a consequence, $\mathbf{x}$ keeps within the predefined tube (25).
-As $L(t)\leq V(t)$, it follows from (25) and (27) that $\|\mathbf{x}_{e}(t)\|\leq\rho\sqrt{1-e^{-2V(t)}}$.
-In addition, $\tilde{d}^{2}/(2\eta)\leq V(t)$ leads to $|\tilde{d}(t)|\leq\sqrt{2\eta V(t)}$.
-As shown in (44), $V(t)$ is eventually bounded by $c/a$, whereby one can get
-$\mathop{\lim\sup}\limits_{t\to\infty}\|\mathbf{x}_{e}(t)\|\leq\rho\sqrt{1-e^{-2c/a}},~{}\mathop{\lim\sup}\limits_{t\to\infty}|\tilde{d}(t)|\leq\sqrt{2\eta c/a}.$
+$L^{\prime}(\bar{\xi}(s))$ $\leq-k_{1}L(\bar{\xi}(s))+\dfrac{d}{4k_{2}}\dfrac{T_{f}-\eta_{f}(s)}{T_{f}}$
+$=-k_{1}L(\bar{\xi}(s))+\dfrac{d}{4k_{2}}e^{-\frac{s}{T_{f}}},$ (34)
+on the set $10\leq\bar{\xi}<$ 1. Integrating both sides of (III-C) leads to
+$L(\bar{\xi}(s))$ $\leq L(\bar{\xi}(0))e^{-k_{1}s}+De^{-\frac{s}{T_{f}}},$ (35)
+where $D:=T_{f}d/(4k_{2}(k_{1}T_{f}-1))>0$ is a constant.
+From (35), it is clear that $L(\bar{\xi})$ is bounded, indicating that $10\leq\bar{\xi}(s)<$ 1 for all $s\geq 0$.
+As a result, $\|\bar{\mathbf{x}}_{e}(s)\|<\rho$ holds for all $s\geq 0$.
+Since $t\to T_{f}$ as $s\to\infty$, it follows from (30a) that $\mathbf{x}_{e}(t)$ keeps within the safe tube (22) for all $t\in[0,T_{f}^{\ast})$.
+In the following, we show that $\mathbf{x}_{e}(t)$ converges to a small neighborhood of origin at $t=T_{f}^{\ast}$.
+It follows from (35) that $\lim_{s\to s^{\ast}}L(\bar{\xi}(s))\leq L^{\ast}:=L(\bar{\xi}(0))e^{-k_{1}s^{\ast}}+De^{-s^{\ast}/T_{f}}$.
+As per the definition of $\eta_{f}(s)$, one can easily deduce that $s\to s^{\ast}:=-T_{f}\ln(\varsigma_{f}/T_{f})$ as $t\to T_{f}^{\ast}$.
+Since $0<\varsigma_{f}\ll T_{f}$, $s^{\ast}$ is a very large value, which renders $L^{\ast}$ very small.
+By (23) and (32), we get $\lim_{s\to s^{\ast}}\|\bar{\mathbf{x}}_{e}(s)\|\leq\rho\sqrt{1-e^{-2L^{\ast}}}$, which implies that $\bar{\mathbf{x}}_{e}$ converges to a small neighborhood of origin as $s\to s^{\ast}$.
+Since $t\to T_{f}^{\ast}$ as $s\to s^{\ast}$ and $\bar{\mathbf{x}}_{e}(s)=\mathbf{x}_{e}(\eta_{f}(s))$ (see (30a)), it can be claimed that $\lim_{t\to T_{f}^{\ast}}\|\mathbf{x}_{e}(t)\|\leq\rho\sqrt{1-e^{-2L^{\ast}}}$.
+2) Consider the time interval $t\in[T_{f}^{\ast},\infty)$, in which $\alpha_{f}^{\ast}(t)=\alpha_{f}(T_{f}^{\ast})\equiv T_{f}/\varsigma_{f}$ and the closed-loop system becomes
+$\dot{\mathbf{x}}_{e}=-k_{1}\alpha_{f}(T_{f}^{\ast})\mathbf{x}_{e}-k_{2}\mathbf{z}+\mathbf{R}(\theta)\mathbf{u}_{d}.$ (36)
+We likewise consider the barrier function $L$ in (32). Taking its derivative w.r.t. $t$ along (36) gets
+$\dot{L}(\xi(t))\leq-\bar{k}L(\xi(t))+\bar{d},$ (37)
+where $\bar{k}:=k_{1}T_{f}/\varsigma_{f}$ and $\bar{d}:=d/(4k_{2})$ are defined for notational concision. Solving (37), we get
+$L(\xi(t))\leq L(\xi(T_{f}^{\ast}))e^{-\bar{k}(t-T_{f}^{\ast})}+\bar{d}/\bar{k},~{}\forall t\geq T_{f}^{\ast},$ (38)
+indicating that $L(\xi(t))$ is bounded on $t\in[T_{f}^{\ast},\infty)$.
+Thus, $\mathbf{x}_{e}(t)$ satisfies the tube constraint (22) for all $t\geq T_{f}^{\ast}$.
+In addition, $L$ remains within the set $L(\xi(t))\leq L^{\ast}+\bar{d}/\bar{k}$ for all $t\geq T_{f}^{\ast}$.
+As a result, the tracking error $\mathbf{x}_{e}$ remains within the residual set $\|\mathbf{x}_{e}(t)\|\leq\rho\sqrt{1-e^{-2(L^{\ast}+\bar{d}/\bar{k})}}$ for all $t\geq T_{f}^{\ast}$.
 […]
 Remark 5.
-Based on the facts that $\|\mathbf{R}^{-1}(\theta)\|=1/\ell$ ($\ell<1$), $\|\bm{\tau}_{d}\|\leq\|\bm{\kappa}_{0}(\mathbf{x}_{d})\|\leq\alpha$ (noting (13), (14), and $\|\bm{\Pi}(\mathbf{x}_{d})\|=1$), and $\|\bm{\varpi}\|\leq\hat{d}\leq d_{m}+\sigma$ (noting (34)), it can be easily verified that the adaptive controller (29) is bounded and its bound can be explicitly expressed as
-$\|\mathbf{u}\|\leq\dfrac{1}{\ell}(k\rho+\alpha+d_{m}+\delta).$ (45)
-This implies that the input constraint $\|\mathbf{u}\|\leq u_{m}$ can be met by judiciously choosing the design parameters such that $k\rho+\alpha+d_{m}+\delta\leq\ell u_{m}$.
-However, it is noted that achieving smaller convergence bounds for $\mathbf{x}_{e}$ and $\tilde{d}$ requires a larger value of $k$.
-Thus, a compromise should be reached according to the robot actuation capability and performance requirement.
-Remark 5.
-As defined in (1), the kinematics of $\theta$ is
-$\dot{\theta}=\omega+u_{d2},$ (46)
-[…]
+Under the InPTC framework, if the radius of the tube is set less than the safety margin $\epsilon$, whilst $T_{f}$ is set no larger than $T$, then the actual trajectory of $\mathbf{x}$ will converge to a small neighborhood of the goal position $\mathbf{x}^{\ast}$ (if no local minima occurs) within the prescribed time $T$, while avoiding collision with all obstacles $\mathcal{O}_{i}$, $i\in\mathbb{I}$ along the way.
+Therefore, the proposed InPTC can achieve prescribed-time collision-free navigation of WMRs.
+Furthermore, inspecting (35) and (38) reveals that choosing a smaller $\varsigma_{f}$ together with larger values for $k_{1}$ and $k_{2}$ decreases the size of residual set.
+However, this comes at the price of higher control gains (see (26)), particularly when selecting a very small $\varsigma_{f}$.
+This, in turn, may excite the unmodeled high frequency dynamics, leading to instability of the closed system.
+Thus, $k_{1}$, $k_{2}$, and $\varsigma_{f}$ should be judiciously chosen to strike a balance between tracking accuracy and system stability.
 IV Simulation Results
-In this section, we demonstrate the proposed control method in a $6.4\,\text{m}\times 3.4\,\text{m}$ rectangular workspace $\mathcal{W}^{\ast}$ cluttered with 8 circular obstacles, denoted as $\mathcal{O}_{i}^{\ast}$, whose centers and radii are listed in Table I.
-The distance between the virtual and actual control points is $\ell=0.05\,\rm m$, the radius of the circumscribed circle around $\bm{P}$ is $r=0.2\,\rm m$, and the clearance constant in Assumption 2 is $h=0.2\,\rm m$.
-The size of the safety margin is set to $\epsilon=0.1\,\rm m$, and the obstacle influence region is specified by $\epsilon^{\ast}=0.2\,\rm m$.
-The goal location is taken as $\mathbf{x}^{\ast}=[2.5,1]^{\top}\,\rm m$.
-Assume that the control input constraint is $u_{m}=1.5$, and the disturbance $\mathbf{u}_{d}$ is of the following form
-$\mathbf{u}_{d}=0.01\times\left[\begin{matrix}\sin(0.2t)+1\\ \cos(0.3t)-2\end{matrix}\right].$ (47)
-According to Remark 4, the planner parameters are chosen as $\alpha=0.03$ and $\beta=0.005$, while the controller parameters are chosen as $\rho=0.06$, $k=0.1$, $\psi=0.005$, $\eta=0.1$, $\gamma=0.01$, $d_{m}=0.03$, $\delta=0.005$, and $\hat{d}(0)=0.01$.
-The simulation duration is $500\,\rm s$, and the sample step is $0.01\,\rm s$.
-We begin by verifying the efficiency of the proposed tangent cone-based path planning method (denoted as “TC-based planner”) in the absence of disturbances.
-For comparison purposes, the PF-based path planning method presented in [30] (denoted as “PF-based planner”) is also simulated.
-A slight modification is made for [30] to facilitate its implementation.
-The attractive and repulsive potentials are chosen as
-$U_{att}(\mathbf{x}_{d})=$ $\dfrac{1}{2}k_{a}\|\mathbf{x}_{d}-\mathbf{x}^{\ast}\|^{2},$
-$U_{rep}(\mathbf{x}_{d})=$ $\dfrac{1}{2}k_{r}\left(\sum_{i=0}^{8}\dfrac{1}{\rho_{i}}\right)\|\mathbf{x}_{d}-\mathbf{x}^{\ast}\|^{2},$
-where $k_{a}$ and $k_{r}$ are positive weights, $\rho_{0}:=1-(x_{d}/2.9)^{20}-(y_{d}/1.4)^{20}$, while $\rho_{i}:=\|\mathbf{x}_{d}-\mathbf{c}_{i}\|^{2}-(r+r_{i}+\epsilon)^{2}$, $i\in\mathbb{I}$.
-The total velocity is then given by
-$\bm{\tau}_{d}=-\nabla U_{att}(\mathbf{x}_{d})-\nabla U_{rep}(\mathbf{x}_{d}).$ (48)
-For fair comparison, we chose the parameters of the PF-based path planner to be $k_{a}=0.05$ and $k_{r}=0.0001$, such that the two planners have nearly identical convergence times.
-Figure 6 shows the obtained robot navigation trajectories of both path planners starting at a set of initial positions (purple points).
-As can be seen, both planners can successfully guide the virtual control point $\bm{P}$ from different initial positions to the goal (red point), while avoiding all obstacle regions $\mathcal{O}_{i}^{\epsilon}$, $i=1,2,...,8$.
-However, the proposed planner yields shorter trajectories than the PF-based one.
-The comparison results of the velocity norm $\|\bm{\tau}_{d}\|$ are depicted in Fig.
-7, from which it is clear that the PF-based planner requires larger velocity than the proposed one.
-Actually, $\bm{\tau}_{d}$ under the proposed planner is upper bounded by $\alpha=0.03$, which is consistent with our analysis in Remark 3.
-The above results demonstrate the effectiveness and superiority of the proposed path planner.
+In this section, we demonstrate the proposed InPTC scheme in a $6.4\,\text{m}\times 3.4\,\text{m}$ rectangular workspace $\mathcal{W}^{\ast}$ cluttered with 8 circular obstacles, denoted as $\mathcal{O}_{i}^{\ast}$, whose centers and radii are listed in Table I.
+The distance between the virtual and actual control points is $\ell=0.05\,\rm m$, the radius of the circumscribed circle around $\bm{P}$ is $r=0.2\,\rm m$.
+The safety margin and influence region of obstacles are set to $\epsilon=0.1\,\rm m$ and $\epsilon^{\ast}=0.2\,\rm m$, respectively.
+The desired goal is $\mathbf{x}^{\ast}=[2.5,1]^{\top}\,\rm m$, and the task completion time is $200\,\rm s$.
+The disturbance $\mathbf{u}_{d}$ is modeled as $\mathbf{u}_{d}=0.01[\sin(0.2t)+1,\cos(0.3t)-2]^{\top}$.
+The simulation duration is $1000\,\rm s$, and the sample step is $0.05\,\rm s$.
 TABLE I: Geometrical details of the obstacles.
@@
 $\mathcal{O}_{4}^{\ast}$ $[-2.1,0.6]^{\top}$ $0.15$ $\mathcal{O}_{8}^{\ast}$ $[1.8,0.7]^{\top}$ $0.15$
-Figure 6: Navigation trajectories of the two path planners.
-Figure 7: Velocity norms of the two path planners.
-We next illustrate the performance of the proposed adaptive tube-following controller (29) (denoted as “TF controller”) in the presence of disturbance given by (47).
-As a case study, the trajectory starting from the 3nd initial position is selected as the reference trajectory (see Fig.
-6).
-For comparison purposes, the extensively used proportional-integral (PI) controller of the following form is simulated:
-$\mathbf{u}=-k_{p}\mathbf{x}_{e}-k_{i}\int_{0}^{t}\mathbf{x}_{e}(\tau)\text{d}\tau,$
-where $k_{p}=0.5$ and $k_{i}=0.3$ are constant gains.
-Furthermore, in order to verify the robustness of the PF-based planner (48) (denoted here as “PF controller”), we execute it in the presence of disturbances and define the difference between the ideal and actual trajectories as the position tracking error $\mathbf{x}_{e}$.
-The closed-loop responses of these three controllers are shown in Fig.
+We begin by illustrating the effectiveness of the proposed prescribed-time planner (19) (denoted as “PTP”).
+For comparison purposes, the APF-based planner presented in [32] (denoted as “APF”) and the CBF-based planner presented in [9] (denoted as “CBF”) are also simulated.
+The attractive and repulsive potentials of the APF planner are designed as $U_{att}(\mathbf{x}_{d})=\frac{k_{a}}{2}\|\mathbf{x}_{d}-\mathbf{x}^{\ast}\|^{2}$, $U_{rep}(\mathbf{x}_{d})=k_{r}\sum_{i=1}^{8}\Upsilon(\text{d}_{\mathcal{O}_{i}}(\mathbf{x}_{d}))$, where $k_{a}$ and $k_{r}$ are positive weights, and $\Upsilon$ is a smooth repulsive function defined as $\Upsilon(z)=(\epsilon^{\ast}-z)^{2}\text{ln}(z-\epsilon)/(z-\epsilon)$ if $z\in(\epsilon,\epsilon^{\ast}]$, and $\Upsilon(z)=0$, if $z>\epsilon^{\ast}$.
+Then, the vector field generated by the APF planner is
+$\bm{\tau}_{d}=-\nabla U_{att}(\mathbf{x}_{d})-\nabla U_{rep}(\mathbf{x}_{d}).$ (39)
+We consider a CBF $f(\mathbf{x}_{d}):=\min_{i\in\{0\}\cup\mathbb{I}}f_{i}(\mathbf{x}_{d})$ together with the safe set :=$\mathcal{F}:=\{\mathbf{x}_{d}\in\mathbb{R}^{2}\mid f(\mathbf{x}_{d})\geq 0\}$, where $f_{0}(\mathbf{x}_{d}):=1-(x_{d}/2.9)^{20}-(y_{d}/1.4)^{20}$ (for workspace boundaries), and $f_{i}(\mathbf{x}_{d}):=\|\mathbf{x}_{d}-\mathbf{c}_{i}\|^{2}-(r+r_{i}+\epsilon)^{2}$, $i\in\mathbb{I}$.
+The CBF planner is obtained by quadratic programming
+$\mathop{\rm argmin}\limits_{\bm{\tau}_{d}\in\mathbb{R}^{2}}$ $~{}\|\bm{\tau}_{d}-\bm{\tau}_{d,\rm des}(\mathbf{x}_{d})\|^{2}$ (CBF-QP)
+s.t. $~{}\nabla f(\mathbf{x}_{d})^{\top}\bm{\tau}_{d}\geq-\gamma f(\mathbf{x}_{d}),$
+where $\alpha>0$ is a user-defined constant. The CBF-QP has an explicit solution of the form
+$\bm{\tau}_{d}=\bm{\tau}_{d,\rm des}(\mathbf{x}_{d})+\bm{\tau}_{d,\rm safe}(\mathbf{x}_{d}),$ (40)
+where $\bm{\tau}_{d,\rm des}(\mathbf{x}_{d}):=\bm{\kappa}_{0}(\mathbf{x}_{d})$, and $\bm{\tau}_{d,\rm safe}$ is determined by
+$\bm{\tau}_{d,\rm safe}(\mathbf{x}_{d}):=\left\{\begin{aligned} &-\dfrac{\nabla f(\mathbf{x}_{d})}{\nabla f(\mathbf{x}_{d})^{\top}\nabla f(\mathbf{x}_{d})}\Psi(\mathbf{x}_{d}),&~{}&\text{if}~{}\Psi<0,\\ &0,&~{}&\text{if}~{}\Psi\geq 0,\end{aligned}\right.$
+with $\Psi(\mathbf{x}_{d}):=\nabla f(\mathbf{x}_{d})^{\top}\bm{\tau}_{d,\rm des}(\mathbf{x}_{d})+\gamma f(\mathbf{x}_{d})$.
+As dictated by [9], the CBF planner (40) ensures that the set $\mathcal{F}$ is forward invariant.
+The APF planner parameters are chosen as $k_{a}=k_{0}$ and $k_{r}=0.1$, while the CBF planner parameter is chosen as $\gamma=0.1$.
+It is important to note that these three planners have the same motion-to-goal control law.
+Their design parameters are listed in Table II.
+TABLE II: Planner and controller parameters.
+Method Parameter
+PTP in (19) $k_{0}=0.01$, $T=200$, $\varsigma=0.5$
+TFC in (26) $\rho=0.06$, $k_{1}=0.8$, $k_{2}=0.001$, $T_{f}=200$, $\varsigma_{f}=3$
+APF in [32] $k_{a}=k_{0}=0.01$, $k_{r}=0.1$
+CBF in [9] $\gamma=0.1$
+The planned trajectories starting at a set of initial positions (purple points) are plotted in Fig.
+6, where the dense arrows denote the magnitude and direction of the resulting vector field from the prescribed-time planner (19) at various points.
+These vector arrows are directly obtained as per (19).
+As shown in Fig.
+6, all three planners successfully guide the WMR from different initial positions to the goal (red point), while avoiding all obstacle regions.
+The comparison results of $\|\mathbf{x}_{d}\|$ and $\|\bm{\tau}_{d}\|$ are depicted in Fig.
+7, from which we find that both the APF and CBF converge slower than the proposed PTP; moreover, their convergence times become longer as the path is longer.
+However, the PTP always converges at $T=200\,\rm s$, regardless of initial positions.
+We further provide quantitative comparison results in Table III, where the 3rd initial position is considered as a case study.
+Although all the three planners offer analytical solutions and have an almost identical time consumption (average time of 100 runs), the convergence time differs greatly.
+The PTP converges at the prescribed time $T=200\,\rm s$, while the APF and CBF take significantly longer, converging in $950\,\rm s$.
+The path length metric also differs, with CBF covering a shorter path compared to the other two methods.
+The maximum and standard deviation (std) of the velocity norm provide additional insights, with PTP exhibiting higher values than APF and CBF in both measures, due to its faster convergence rate.
+Figure 6: Planned trajectories of the three path planners.
+(a) PTP in our work
+(b) APF in [32]
+(c) CBF in [9]
+Figure 7: Control responses of different path planners.
+TABLE III: Planning performance comparison.
+Method Time consumption Convergence time Path length Maximum of $\|\bm{\tau}_{d}\|$ std of $\|\bm{\tau}_{d}\|$ Obstacle avoidance Prescribed time
+PTP in (19) 0.3136 s 200 s 6.4411 m 0.0548 0.0148
+APF in [32] 0.3196 s 950 s 6.7394 m 0.0548 0.0096
+CBF in [9] 0.3474 s 950 s 6.3976 m 0.0479 0.0102
+In the following, we show the effectiveness and performance of the proposed tube-following controller (26) (denoted as “TFC”) with parameters given in Table II.
+As a case study, the planned trajectory starting from the 3rd initial position is selected as the reference trajectory.
+To verify the robustness of the APF and CBF planners, we execute them in the presence of disturbances and define the difference between the reference and actual trajectories as the position tracking error $\mathbf{x}_{e}$.
+The closed-loop tracking responses of three controllers are shown in Fig.
 8.
-In the top left subfigure of Fig.
-8, the tracking error $\mathbf{x}_{e}$ under both the TF and PI controllers is shown to remain within a small neighborhood of the origin (but quantitatively, the TF controller has higher tracking accuracy than the PI controller), while satisfying the tube constraint (24).
-However, the PF controller yields a larger tracking error and violates the tube constraint.
-The latter may result in collision with obstacles or the workspace boundary.
-The WMR’s heading angle is plotted in the top right subfigure of Fig.
-8, from which it is clear that $\theta$ under the TF and PI controllers is bounded and converges to a nearly constant value (precisely, the TF controller has a higher stability), while under the PF controller $\theta$ continuously increases.
-This implies that the PF controller (i.e., path planner (48)) is susceptible to disturbances.
-The control input norm $\|\mathbf{u}\|$ is shown in the bottom left subfigure of Fig.
-8.
-Obviously, the control input $\mathbf{u}$ of the TF and PI controllers satisfies the magnitude limits $\|\mathbf{u}\|\leq u_{m}$, while the PF controller fails.
-The parameter estimate $\hat{d}$ is given in the bottom right subfigure of Fig.
-8.
-As can be seen, $\hat{d}$ remains within the set $\Omega_{\delta}$, indicating the effectiveness of the projection adaptive law (31).
-Finally, the position tracking trajectory of the proposed TF controller (29) is shown in Fig.
+In Fig.
+8(a), the tracking error $\mathbf{x}_{e}$ under the proposed TFC is shown to converge to a residual set of $3.74\times 10^{-4}\,\rm m$ after the prescribed time $T_{f}=200\,\rm s$, while complying with the tube constraint defined by (22).
+In contrast, the APF and CBF controllers exhibit larger stead-state errors of $0.43\,\rm m$ and $0.32\,\rm m$, respectively, whilst violate the tube constraint, leading to an increased risk of collision.
+The WMR’s heading angle is provided in Fig.
+8(b), from which it is clear that $\theta$ under the TFC is bounded and converges to a nearly constant value, while under the APF and CBF, $\theta$ continuously increases.
+This implies that both the APF and CBF controllers are susceptible to disturbances.
+The control norm $\|\mathbf{u}\|$ is shown in Fig.
+8(c).
+To further show the prescribed-time convergence property of the proposed TFC, the tracking error norm under different prescribed time $T_{f}$ is given in Fig.
+8(d).
+As illustrated, the proposed controller can precisely and flexibly set the convergence time in advance.
+The position tracking trajectory of the proposed TFC is shown in Fig.
 9.
-Intuitively, the virtual control point $\bm{P}$ accurately tracks the reference trajectory, while remaining with the predefined tube.
-Therefore, the WMR reaches the goal position without colliding with any obstacles (dark gray balls), as clearly seen in Fig.
-9.
-Overall, the proposed integrated path planning and tube-following control scheme can successfully drive the WMR to reach the predefined goal without colliding with any obstacles, despite the presence of input disturbances.
-Figure 8: Closed-loop control responses.
+Intuitively, $\mathbf{x}$ accurately tracks the reference trajectory, while remaining within the safe tube.
+This implies that, under the proposed InPTC, the WMR reaches the goal within the prescribed time $T=200\,\rm s$, while avoiding collision with any obstacles (dark gray balls).
+Figure 8: Comparisons of the closed-loop tracking responses.
 Figure 9: Position tracking trajectory with WMR snapshots.
 V Experiments
-In this section, several experiments are carried out to verify the effectiveness of the proposed method.
-V-A Experimental setup
-In the experiments, several simple scenarios are considered where the WMR is tasked with navigating towards a designated goal location, while ensuring avoidance of any obstacles en route.
 The experimental platform is shown in Fig.
 10, which consists of a Mona robot [33] with nonholonomic dynamics, several hot obstacles containing iron powders that can heat up themselves, a RGB camera to observe the pose and location of the WMR, a thermal camera to obtain the size and location of the hot obstacles, and a control PC to collect feedback data and send motion commands.
-The Wi-Fi communication (at a rate of $10\,\rm Hz$) between the control PC and the robot is built by ROS.
-All experiments are performed in an arena with a size of $\rm 1.8\,m\times 0.8\,m$.
-It should be emphasized that since the WMR in our lab has a very limited sensing capability, the distance between the WMR and obstacle boundary is obtained in real-time from the RGB and thermal cameras, rather than sensors onboard the WMR.
-Regarding the robot configuration, we set the radius of the circle surrounding the robot to $r=0.06\,\text{m}$, the parameter for the change of coordinates to $\ell=-0.02$, the safety margin to $\epsilon=0.04\,\text{m}$, and the obstacle influence region to $\epsilon^{*}=0.05\,\text{m}$.
-In all experiments, the parameters of the path planner are set to $\alpha=0.03$ and $\beta=0.005$, while the controller parameters are set to $\rho=0.03$, $k=1.5$, $\psi=0.005$, $\eta=0.1$, $\gamma=0.01$, $d_{m}=0.036$, $\delta=0.005$, and $d(0)=0.01$.
-V-B Experimental results
-We here report the results of two representative experiments, where the extracted thermal image and the RGB image are merged to record the experimental process.
-In the first experiment, six obstacles are placed in the arena.
-The WMR starts from the left side of the arena (coordinates: $[0.27,0.21]^{\top}\text{m}$), and is tasked with reaching the goal location situated on the right side (coordinates: $[1.73,0.54]^{\top}\text{m}$).
+The Wi-Fi communication (at a rate of $10\,\rm Hz$) between the control PC and robot is built by ROS.
+All experiments are performed in a $\rm 2.78\,m\times 1.4\,m$ arena.
+Since the Mona robot has a very limited onboard sensing capability, the distance between the robot and obstacle boundary is obtained in real-time from the RGB and thermal cameras.
+The radius of the circle surrounding the robot is set to $r=0.06\,\text{m}$, the parameter for the change of coordinates is $\ell=-0.02\,\rm m$.
+The design parameters of the InPTC are listed in Table IV.
+In experiments, the disturbances could come from measurement noises, communication delay, imperfect wheel alignment, motor or actuator asymmetry, etc.
+In the following, we report the results of two representative experiments, where the extracted thermal image and the RGB image are merged to record the experimental process.
+Figure 10: The experiment setup: (a) platform; (b) cameras; (c) Mona robot.
+TABLE IV: Design parameters of the InPTC.
+Method Parameter
+Planner in (19) $k_{0}=0.01$, $T=250$, $\varsigma=0.5$
+Controller in (26) $\rho=0.06$, $k_{1}=0.01$, $k_{2}=0.003$, $T_{f}=150$, $\varsigma_{f}=2$
+In the first experiment, eight circular obstacles are placed in the arena, and the safety margin and influence region of obstacles are set to $\epsilon=0.04\,\text{m}$ and $\epsilon^{*}=0.05\,\text{m}$, respectively.
+The WMR starts from the right side of the arena (coordinates: $[2.45,1.27]^{\top}\text{m}$), and is tasked with reaching the goal located on the right side (coordinates: $[0.44,0.35]^{\top}\text{m}$).
+The task is required to be completed within $250\,\rm s$.
 The results are shown in Fig.
-11, from which it can be observed that the proposed path planner generates a collision-free navigation path for the WMR (i.e., the red line in Fig.
-11(a)), and the tube-following controller guarantees that the WMR tracks the desired path along with the predefined safe tube.
-The latter can be seen from Fig.
-11(c), where the norm of the position tracking error is smaller than the tube width $\rho=0.03\text{m}$ all the time.
-In the second experiment, we add two more hot obstacles to the arena, which makes the environment more complex for the WMR.
-In this case, the WMR starts from the right side of the arena (coordinate: $[1.83,0.68]^{\top}\text{m}$) and is required to reach a goal location at the left side (coordinate: $[0.20,0.40]^{\top}\text{m}$).
-The experiment results are shown in Fig.
-12, where we can observe similar outcomes as in the first experiment.
-Our path planner successfully generates a collision-free navigation path for the WMR and the tube-following controller drives the WMR to track the generated path without running out the predefined tube (see Fig.
+11, from which it is clear that the proposed planner in (19) generates a smooth, collision-free reference trajectory that converges to the goal at the preassigned time $T=250\,\rm s$ (i.e., the red line in Fig.
+11(a)).
+While the proposed controller in (26) achieves trajectory tracking along with the safe tube (see Fig.
+11(c)), with a tracking accuracy higher than $4.9\times 10^{-3}\,\rm m$ after the preassigned time $T_{f}=150\,\rm s$ (see Fig.
+11(d)).
+As seen in Fig.
+11(e), the robot’s heading angle remains nearly constant after the task completion time $T=250\,\rm s$.
+The above results demonstrate that the proposed InPTC scheme enables the WMR to safely accomplish the navigation task within the required time of $250\,\rm s$.
+In the second experiment, we consider a more complex environment with six obstacles of various shapes.
+The safety margin and influence region of obstacles are set to $\epsilon=0.08\,\text{m}$ and $\epsilon^{*}=0.1\,\text{m}$, respectively.
+The WMR starts from the left side of the arena (coordinate: $[0.3,0.5]^{\top}\text{m}$) and is required to reach a goal location at the right side (coordinate: $[2.5,1.0]^{\top}\text{m}$).
+The results are depicted in Fig.
+12, hich shows outcomes analogous to those observed in the initial experiment.
+In this case, the proposed InPTC still generates a smooth and safe reference trajectory with prescribed-time convergence (see Fig.
 12(c)).
-Furthermore, we perform multiple trials with different initial and goal locations to further test the proposed method.
-From these experiments, we consistently observe the effectiveness of our proposed robot navigation algorithm.
-More details can be found in the supplementary video (https://vimeo.com/895801720).
-Figure 10: The experiment setup: (a) experimental platform; (b) cameras; (c) Mona robot.
-Figure 11: The experimental scenario with six obstacles: (a) the generated path (red line) and the actual trajectory (black line) of the WMR; (b) screenshots of the experiment; (c) position error; (d) WMR’s orientation; (e) parameter estimation $\hat{d}$.
-Figure 12: The experimental scenario with eight obstacles: (a) the generated path (red line) and the actual trajectory (black line) of the WMR; (b) screenshots of the experiment; (c) position error; (d) WMR’s orientation; (e) parameter estimation $\hat{d}$.
+It also achieves prescribed-time tube-following control with an accuracy of $4.7\times 10^{-3}\,\rm m$ (see Fig.
+12(d)), thus enabling the WMR to accomplish the navigation task within the prescribed time $T=250\,\rm s$ (see Fig.
+12(a)).
+Furthermore, we perform multiple trials with different initial and goal locations to further test the proposed InPTC.
+From these experiments, we consistently observe the effectiveness of InPTC in achieving tracking accuracy higher than $0.01\,\rm m$ and task completion time less than $250\,\rm s$.
+More details are provided in the supplementary video (https://vimeo.com/895801720).
+Figure 11: The experiment with circular obstacles: (a) WMR’s motion trajectory; (b) screenshots of the experiment; (c) reference position norm; (d) position error norm; (e) WMR’s heading angle.
+Figure 12: The experiment with circular and polygonal obstacles: (a) WMR’s motion trajectory; (b) screenshots of the experiment; (c) reference position norm; (d) position error norm; (e) WMR’s heading angle.
 VI Conclusion
-In this paper, we have proposed an integrated path planning and tube-following control scheme for collision-free navigation of WMRs in a convex workspace cluttered with multiple spherical obstacles.
-A path planner was designed using Bouligand’s tangent cones to generate safe, collision-free motion paths for the WMR.
-The resulting vector field asymptotically guides the WMR from almost all initial configurations in the free space towards the goal location, without colliding with any obstacles along the way.
-Then, an adaptive tube-following controller was derived, which guarantees that the actual trajectory of the WMR remains within a predefined safe tube around the reference trajectory, despite the presence of disturbances.
-The tube-following control significantly enhances the safety of robot navigation in the perturbed scenarios.
-In particular, the proposed closed-loop navigation strategy can be implemented using only the local sensing information.
-Finally, numerical simulations and experiments illustrate the effectiveness of the proposed autonomous navigation algorithm.
-Future work will focus on reactive collision-free navigation of multiple WMRs.
-[Proof of Theorem 2] From (18) and (19), it follows that $\bm{\tau}(\mathbf{x})\in\mathbf{T}_{\mathcal{X}_{\epsilon}}(\mathbf{x})$ for all $\mathbf{x}\in\mathcal{X}_{\epsilon}$, which according to Theorem 1 suggests that the free space $\mathcal{X}_{\epsilon}$ is forward invariant.
-As previously stated, $\bm{\tau}(\mathbf{x})$ is piecewise continuously differentiable, indicating that it is locally Lipschitz on its domain $\mathcal{X}$ [32].
+In this article, we propose a prescribed-time path planning algorithm for wheeled mobile robots (WMRs) to generate a collision-free reference trajectory that can converge to the goal location at a prescribed finite time.
+The derived planner is then paired with a tube-following controller that allows WMRs to achieve high-precision trajectory tracking within a prescribed time, while remaining within a predefined “safe tube” around the reference trajectory.
+The resulting integrated planning and tube-following control (InPTC) scheme achieves prescribed-time collision-free navigation of WMRs, despite the presence of disturbances.
+The salient features of the proposed InPTC are two-fold: 1) it enables WMRs to accomplish safe navigation within a finite time that can be flexibly preassigned according to the mission requirements; and 2) it enhances the navigation safety of WMRs in practical implementations.
+Simulation and experimental results show the effectiveness of InPTC.
+Future work will focus on extending the proposed method to achieve prescribed-time safe navigation of WMRs in environments with dynamic or intersecting obstacles.
+[Proof of Theorem 2]
+From (11) and (12), it follows that $\mathbf{h}(\mathbf{x})\in\mathbf{T}_{\mathcal{X}_{\epsilon}}(\mathbf{x})$ for all $\mathbf{x}\in\mathcal{X}_{\epsilon}$, which according to Theorem 1 suggests that the free space $\mathcal{X}_{\epsilon}$ is forward invariant.
+As previously stated, $\mathbf{h}(\mathbf{x})$ is piecewise continuously differentiable, indicating that it is locally Lipschitz on its domain $\mathcal{X}$ [34].
 Moreover, the free space $\mathcal{X}{\epsilon}$, as a compact subset of $\mathcal{X}$, is forward invariant.
-Then, it follows from [33, Theorem 3.3] that the closed-loop kinematics (11) with $\mathbf{x}(0)\in\mathcal{X}_{\epsilon}$ has a unique solution that is defined for all $t\geq 0$.
-Subsequently, we show that the unique solution converges to a set of stationary points.
+Then, it follows from [35, Theorem 3.3] that the closed-loop kinematics (7) under $\bm{\tau}=\mathbf{h}(\mathbf{x})$ with $\mathbf{x}(0)\in\mathcal{X}_{\epsilon}$ has a unique solution that is defined for all $t\geq 0$.
+We next show that the unique solution converges to a set of stationary points.
 To this end, we consider the following LFC:
@@
 For ease of analysis, (11) is rewritten as
-$\bm{\tau}(\mathbf{x})=\underbrace{(c(\mathbf{x})\mathbf{I}_{2}+(1-c(\mathbf{x}))\mathbf{\Pi}(\mathbf{x}))}_{\mathbf{\Theta}(\mathbf{x})}\bm{\kappa}_{0}(\mathbf{x}),$ (50)
+$\mathbf{h}(\mathbf{x})=\underbrace{(c(\mathbf{x})\mathbf{I}_{2}+(1-c(\mathbf{x}))\mathbf{\Pi}(\mathbf{x}))}_{\mathbf{\Theta}(\mathbf{x})}\bm{\kappa}_{0}(\mathbf{x}),$ (42)
 with
@@
 Taking the time derivative of $W(\mathbf{x})$ along (7) and noting (9) and (42), one gets
-$\dot{W}(\mathbf{x})=-k_{0}(\mathbf{x})(\mathbf{x}-\mathbf{x}^{\ast})^{\top}\mathbf{\Theta}(\mathbf{x})(\mathbf{x}-\mathbf{x}^{\ast}).$ (52)
-Since $k_{0}(\mathbf{x})>0$ and $\mathbf{\Pi}(\mathbf{x})$ is positive semidefinite, one has
+$\dot{W}(\mathbf{x})=-k_{0}(\mathbf{x}-\mathbf{x}^{\ast})^{\top}\mathbf{\Theta}(\mathbf{x})(\mathbf{x}-\mathbf{x}^{\ast}).$ (44)
+Since $k_{0}>0$ and $\mathbf{\Pi}(\mathbf{x})$ is positive semidefinite, one has
 $\dot{W}(\mathbf{x})\leq 0,~{}\forall\mathbf{x}\in\mathcal{X}_{\epsilon},$ (45)
@@
 From LaSalle’s invariance principle, it follows that the solution of (7) asymptotically converges to the set $\{\mathbf{x}\in\mathcal{X}_{\epsilon}\mid\dot{W}(\mathbf{x})=0\}$, i.e., a set of stationary points.
-One can infer from (19) and (52) that the stationary points correspond to either $\mathbf{x}=\mathbf{x}^{\ast}$ or the points satisfying $\text{d}_{\mathcal{O}}(\mathbf{x})=\epsilon$ and $\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})/\|\bm{\kappa}_{0}(\mathbf{x})\|=1$.
-Therefore, for any $\mathbf{x}(0)\in\mathcal{X}_{\epsilon}$, the solution of (11) converges to the set $\{\mathbf{x}^{\ast}\}\cup_{i=1}^{n}\mathcal{S}_{i}$.
+One can infer from (12) and (44) that the stationary points correspond to either $\mathbf{x}=\mathbf{x}^{\ast}$ or the points satisfying $\text{d}_{\mathcal{O}}(\mathbf{x})=\epsilon\wedge\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})/\|\bm{\kappa}_{0}(\mathbf{x})\|=1$.
+Therefore, for any $\mathbf{x}(0)\in\mathcal{X}_{\epsilon}$, the solution of (7) converges to the set $\mathcal{C}=\{\mathbf{x}^{\ast}\}\cup_{i=1}^{n}\{\mathbf{s}_{i}\}$.
 The set contains $n+1$ stationary points, all of which are isolated due to Assumption 2.
-Geometrically, the undesired stationary point $\mathbf{s}_{i}\in\mathcal{S}_{i}$ (on the boundary of $\mathcal{O}_{i}^{\epsilon}$) and $\mathbf{x}^{\ast}$ are collinear with $\mathbf{c}_{i}$, but located on the opposite sides of $\mathbf{c}_{i}$, as seen in Fig.
+Geometrically, the undesired stationary point $\mathbf{s}_{i}$ (on the boundary of $\mathcal{O}_{i}^{\epsilon}$) and $\mathbf{x}^{\ast}$ are collinear with $\mathbf{c}_{i}$, but located on the opposite sides of $\mathbf{c}_{i}$, as seen in Fig.
 13.
@@
 The isolated stationary points $\mathbf{s}_{i}$, $i\in\mathbb{I}$ may prevent us from achieving the objective (motion-to-goal).
-To check the stability of $\mathbf{s}_{i}$, we define a small neighborhood of $\mathbf{s}_{i}$, i.e., an open ball $\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$ with $0<\epsilon^{\prime}\leq\min\{\epsilon,\epsilon^{\ast}-\epsilon\}$, such that $\bf{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})>0$ holds for all $\mathbf{x}\in\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$.
+To check the stability of $\mathbf{s}_{i}$, we define a small neighborhood of $\mathbf{s}_{i}$, i.e., an open ball $\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$ with $0<\epsilon^{\prime}\leq\min\{\epsilon,\epsilon^{\ast}-\epsilon\}$, such that $\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})>0$ holds for all $\mathbf{x}\in\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$.
 When restricted on $\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$, the feasible set of initial configurations is $\mathcal{F}_{i}:=\mathcal{X}_{\epsilon}\cap\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$, and $\mathbf{x}$ evolves according to
-$\dot{\mathbf{x}}=-k_{0}(\mathbf{x})(\mathbf{I}_{2}-\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))\mathbf{b}(\mathbf{x})\mathbf{b}^{\top}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast}).$ (55)
+$\dot{\mathbf{x}}=-k_{0}(\mathbf{I}_{2}-\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))\mathbf{b}(\mathbf{x})\mathbf{b}^{\top}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast}).$ (47)
 Further define two sets
@@
 Note that for all $\mathbf{x}\in\mathcal{M}_{i}$, $\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})/\|\bm{\kappa}_{0}(\mathbf{x})\|=1$ holds, whereby (47) reduces to
-$\dot{\mathbf{x}}=-k_{0}(\mathbf{x})(1-\phi(\text{d}_{\mathcal{O}}(\mathbf{x})))(\mathbf{x}-\mathbf{x}^{\ast}).$ (58)
+$\dot{\mathbf{x}}=-k_{0}(1-\phi(\text{d}_{\mathcal{O}}(\mathbf{x})))(\mathbf{x}-\mathbf{x}^{\ast}).$ (50)
 From (13) and (48), it is evident that $0\leq\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))\leq 1$ on the set $\mathcal{M}_{i}$ and $\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))=1$ only when $\mathbf{x}=\mathbf{s}_{i}$.
-Moreover, $k_{0}(\mathbf{x})$ is strictly larger than 00 by construction.
-Hence, for all $\mathbf{x}\in\mathcal{M}_{i}\setminus\{\mathbf{s}_{i}\}$, the control vector $\bm{\tau}(\mathbf{x})$ (i.e., the right-hand side of (58)) points directly towards $\mathbf{s}_{i}$ and becomes zero when $\mathbf{x}=\mathbf{s}_{i}$, indicating that the set $\mathcal{M}_{i}$ is forward invariant.
+Hence, for all $\mathbf{x}\in\mathcal{M}_{i}\setminus\{\mathbf{s}_{i}\}$, the velocity vector $\mathbf{h}(\mathbf{x})$ (i.e., the right-hand side of (50)) points directly towards $\mathbf{s}_{i}$ and becomes zero when $\mathbf{x}=\mathbf{s}_{i}$, indicating that the set $\mathcal{M}_{i}$ is forward invariant.
 Consider the following LFC:
@@
 The time derivative of $W_{i}(\mathbf{x})$ along (50) is given by
-$\dot{W}_{i}(\mathbf{x})=-k_{0}(\mathbf{x})(1-\phi(\text{d}_{\mathcal{O}}(\mathbf{x})))(\mathbf{x}-\mathbf{s}_{i})^{\top}(\mathbf{x}-\mathbf{x}^{\ast}).$ (60)
+$\dot{W}_{i}(\mathbf{x})=-k_{0}(1-\phi(\text{d}_{\mathcal{O}}(\mathbf{x})))(\mathbf{x}-\mathbf{s}_{i})^{\top}(\mathbf{x}-\mathbf{x}^{\ast}).$ (52)
 Since $(\mathbf{x}-\mathbf{s}_{i})^{\top}(\mathbf{x}-\mathbf{x}^{\ast})\geq 0$ for all $\mathbf{x}\in\mathcal{M}_{i}$, one can conclude from (52) that $\dot{W}_{i}(\mathbf{x})\leq 0$ on the set $\mathcal{M}_{i}$ and $\dot{W}_{i}(\mathbf{x})=0$ only occurs at $\mathbf{x}=\mathbf{s}_{i}$.
@@
 As such, (47) becomes
-$\dot{\mathbf{x}}=-k_{0}(\mathbf{x})\mathbf{Q}(\mathbf{b}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast}),$ (61)
+$\dot{\mathbf{x}}=-k_{0}\mathbf{Q}(\mathbf{b}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast}),$ (53)
 where $\mathbf{Q}(\mathbf{b}(\mathbf{x})):=\mathbf{I}_{2}-\mathbf{b}(\mathbf{x})\mathbf{b}^{\top}(\mathbf{x})$ is defined for notational brevity.
@@
 Then, taking the time derivative of $W_{i}$ along (53) and noting (46) and the fact that $(\mathbf{x}-\mathbf{c}_{i})^{\top}\mathbf{Q}(\mathbf{b}(\mathbf{x}))=0$, we get
-$\dot{W}_{i}(\mathbf{x})=$ $~{}k_{0}(\mathbf{x})\alpha_{i}(\mathbf{x}-\mathbf{x}^{\ast})^{\top}\mathbf{Q}(\mathbf{b}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast})$
-$-k_{0}(\mathbf{x})(1-\alpha_{i})(\mathbf{x}-\mathbf{c}_{i})^{\top}\mathbf{Q}(\mathbf{b}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast})$
-$=$ $~{}k_{0}(\mathbf{x})\alpha_{i}(\mathbf{x}-\mathbf{x}^{\ast})^{\top}\mathbf{Q}(\mathbf{b}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast})$
-$=$ $~{}k_{0}(\mathbf{x})\alpha_{i}(1-\cos\psi)\|\mathbf{x}-\mathbf{x}^{\ast}\|^{2}>0,$ (62)
+$\dot{W}_{i}(\mathbf{x})=$ $~{}k_{0}\alpha_{i}(\mathbf{x}-\mathbf{x}^{\ast})^{\top}\mathbf{Q}(\mathbf{b}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast})$
+$-k_{0}(1-\alpha_{i})(\mathbf{x}-\mathbf{c}_{i})^{\top}\mathbf{Q}(\mathbf{b}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast})$
+$=$ $~{}k_{0}\alpha_{i}(\mathbf{x}-\mathbf{x}^{\ast})^{\top}\mathbf{Q}(\mathbf{b}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast})$
+$=$ $~{}k_{0}\alpha_{i}(1-\cos\psi)\|\mathbf{x}-\mathbf{x}^{\ast}\|^{2}>0,$ (54)
 where $\psi$ is the angle between the vectors $\mathbf{b}(\mathbf{x})$ and $\bm{\kappa}_{0}(\mathbf{x})$ and satisfies $20<\psi<\pi/$ 2 for all $\mathbf{x}\in\mathcal{N}_{i}\setminus\{\mathbf{s}_{i}\}$.
@@
 whose time derivative along (53) is given by
-$\dot{V}_{i}(\mathbf{x})=-k_{0}(\mathbf{x})(\mathbf{x}-\mathbf{c}_{i})^{\top}\mathbf{Q}(\mathbf{b}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast})=0,$ (64)
+$\dot{V}_{i}(\mathbf{x})=-k_{0}(\mathbf{x}-\mathbf{c}_{i})^{\top}\mathbf{Q}(\mathbf{b}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast})=0,$ (56)
 showing that for all $\mathbf{x}\in\mathcal{N}_{i}\setminus\{\mathbf{s}_{i}\}$, the control vector is tangent to the boundary of $\mathcal{O}_{i}^{\epsilon}$, thus directing the robot to move along $\partial\mathcal{O}_{i}^{\epsilon}$.
@@
 Actually, for any $\mathbf{x}\in\mathcal{F}_{i}\setminus(\mathcal{M}_{i}\cup\mathcal{N}_{i})$, it holds that $\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))\in(0,1)$, and $\bf{\kappa}_{0}(\mathbf{x})$ and $\mathbf{b}(\mathbf{x})$ are not collinear.
-Moreover, we can decompose $\bm{\tau}(\mathbf{x})=\mathbf{\Pi}(\mathbf{x})\bm{\kappa}{0}(\mathbf{x})$ into two components: the component $\bm{\tau}(\mathbf{x})_{\parallel}=\mathbf{b}(\mathbf{x})\mathbf{b}^{\top}(\mathbf{x})\bm{\tau}(\mathbf{x})$ parallel to $\mathbf{b}(\mathbf{x})$ and the component $\bm{\tau}(\mathbf{x})_{\perp}=(\mathbf{I}_{2}-\mathbf{b}(\mathbf{x})\mathbf{b}^{\top}(\mathbf{x}))\bm{\tau}(\mathbf{x})$ perpendicular to $\mathbf{b}(\mathbf{x})$, as shown in Fig.
+Moreover, we can decompose $\mathbf{h}(\mathbf{x})=\mathbf{\Pi}(\mathbf{x})\bm{\kappa}_{0}(\mathbf{x})$ into two components: the component $\mathbf{h}(\mathbf{x})_{\parallel}=\mathbf{b}(\mathbf{x})\mathbf{b}^{\top}(\mathbf{x})\mathbf{h}(\mathbf{x})$ parallel to $\mathbf{b}(\mathbf{x})$ and the component $\mathbf{h}(\mathbf{x})_{\perp}=(\mathbf{I}_{2}-\mathbf{b}(\mathbf{x})\mathbf{b}^{\top}(\mathbf{x}))\mathbf{h}(\mathbf{x})$ perpendicular to $\mathbf{b}(\mathbf{x})$, as shown in Fig.
 13.
@@
 $=$ $~{}(1-\phi(\text{d}_{\mathcal{O}}(\mathbf{x})))\mathbf{b}^{\top}(\mathbf{x})\bm{\kappa}_{0}(\mathbf{x})>0,$ (57)
-which implies that $\bm{\tau}(\mathbf{x}){\parallel}$ and the bearing vector $\mathbf{b}(\mathbf{x})$ have the same direction, causing $\bm{\tau}(\mathbf{x}){\parallel}$ to point towards $\mathbf{c}_{i}$.
-A straightforward calculation yields $(\mathbf{I}_{2}-\mathbf{b}(\mathbf{x})\mathbf{b}^{\top}(\mathbf{x})\mathbf{\Pi}(\mathbf{x})=\mathbf{Q}(\mathbf{b}(\mathbf{x}))$, which allows us to express $\bm{\tau}(\mathbf{x})_{\perp}=\mathbf{Q}(\mathbf{b}(\mathbf{x}))\bm{\kappa}_{0}(\mathbf{x})$.
-Under $\bm{\tau}(\mathbf{x})_{\perp}$, (55) turns to be (61), which results in $\dot{W}_{i}(\mathbf{x})>0$ as seen in (VI).
-Thus, $\bm{\tau}(\mathbf{x})_{\perp}$ generates a velocity component tangent to $\mathbf{b}$, causing $\mathbf{x}$ to move away from $\mathbf{s}_{i}$.
-Since $\bm{\tau}(\mathbf{x})$ and the vector $\mathbf{s}_{i}-\mathbf{x}$ are situated on opposite sides of $\mathbf{b}(\mathbf{x})$ (see Fig.
+which implies that $\mathbf{h}(\mathbf{x})_{\parallel}$ and the bearing vector $\mathbf{b}(\mathbf{x})$ have the same direction, causing $\mathbf{h}(\mathbf{x})_{\parallel}$ to point towards $\mathbf{c}_{i}$.
+A straightforward calculation yields $(\mathbf{I}_{2}-\mathbf{b}(\mathbf{x})\mathbf{b}^{\top}(\mathbf{x})\mathbf{\Pi}(\mathbf{x})=\mathbf{Q}(\mathbf{b}(\mathbf{x}))$, which allows us to express $\mathbf{h}(\mathbf{x})_{\perp}=\mathbf{Q}(\mathbf{b}(\mathbf{x}))\bm{\kappa}_{0}(\mathbf{x})$.
+Under $\mathbf{h}(\mathbf{x})_{\perp}$, (47) turns to be (53), which results in $\dot{W}_{i}(\mathbf{x})>0$ as seen in (VI).
+Thus, $\mathbf{h}(\mathbf{x})_{\perp}$ generates a velocity component tangent to $\mathbf{b}$, causing $\mathbf{x}$ to move away from $\mathbf{s}_{i}$.
+Since $\mathbf{h}(\mathbf{x})$ and the vector $\mathbf{s}_{i}-\mathbf{x}$ are situated on opposite sides of $\mathbf{b}(\mathbf{x})$ (see Fig.
 13), $\mathbf{x}$ cannot move toward $\mathbf{s}_{i}$ on the set $\mathcal{F}_{i}\setminus(\mathcal{M}_{i}\cup\mathcal{N}_{i})$.
 From Fig.
-13, one can intuitively observe that $\bm{\tau}(\mathbf{x})$ guides $\mathbf{x}$ towards the boundary of $\mathcal{O}_{i}^{\epsilon}$ (attributed to $\bm{\tau}(\mathbf{x})_{\parallel})$, while simultaneously keeping away from the manifold $\mathcal{M}_{i}$ (attributed to $\bm{\tau}(\mathbf{x})_{\perp}$).
+13, one can intuitively observe that $\mathbf{h}(\mathbf{x})$ guides $\mathbf{x}$ towards the boundary of $\mathcal{O}_{i}^{\epsilon}$ (attributed to $\mathbf{h}(\mathbf{x})_{\parallel})$, while simultaneously keeping away from the manifold $\mathcal{M}_{i}$ (attributed to $\mathbf{h}(\mathbf{x})_{\perp}$).
 Therefore, for any $\mathbf{x}(0)\in\mathcal{F}_{i}\setminus(\mathcal{M}_{i}\cup\mathcal{N}_{i})$, the solution of (7) will either directly leaves the ball $\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$, or first converges to the set $\mathcal{N}_{i}\setminus\{\mathbf{s}_{i}\}$ and then leaves the ball $\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$ along $\partial\mathcal{O}_{i}^{\epsilon}$ (as shown in Case 2).
@@
 On the ball $\mathcal{B}(\mathbf{x}^{\ast},r^{\ast})$, the kinematics reduces to
-$\dot{\mathbf{x}}=-k_{0}(\mathbf{x})(\mathbf{x}-\mathbf{x}^{\ast}),$ (66)
-It can be easily verified that $k_{0}(\mathbf{x})\geq\alpha/\sqrt{(r^{\ast})^{2}+\beta^{2}}$ on the ball $\mathcal{B}(\mathbf{x}^{\ast},r^{\ast})$.
-Given this, (66) implies the local exponential stability of $\mathbf{x}=\mathbf{x}^{\ast}$.
-Furthermore, by using the inequality $|z|-z^{2}/\sqrt{z^{2}+\beta^{2}}\leq\beta$, $\forall z\in\mathbb{R}$, (52) reduces to
-$\dot{W}_{i}(\mathbf{x})\leq$ $-\alpha\|\mathbf{x}-\mathbf{x}^{\ast}\|+\alpha\beta$
-$\leq$ $-\sqrt{2}\alpha W_{i}(\mathbf{x})^{\frac{1}{2}}+\alpha\beta,$ (67)
+$\dot{\mathbf{x}}=-k_{0}(\mathbf{x}-\mathbf{x}^{\ast}),$ (58)
 […]